### v1

# Accelerating Ill-conditioned Hankel Matrix RecoveryHanQin Cai, Longxiu Huang, Xiliang Lu, and Juntao You

Abstract
This paper studies the robust Hankel recovery problem, which simultaneously removes the sparse outliers and fulfills missing entries from the partial observation. We propose a novel non-convex algorithm, coined Hankel Structured Newton-Like Descent (HSNLD), to tackle the robust Hankel recovery problem. HSNLD is highly efficient with linear convergence, and its convergence rate is independent of the condition number of the underlying Hankel matrix. The recovery guarantee has been established under some mild conditions. Numerical experiments on both synthetic and real datasets show the superior performance of HSNLD against state-of-the-art algorithms.
keywords:
Ill-conditioned Hankel matrix, robust matrix completion, Newton-like descent
{MSCcodes}
15A29, 15A83, 47B35, 90C17, 90C26, 90C53
1 Introduction
The problem of low-rank Hankel matrix recovery has widely appeared in magnetic resonance imaging[26, 25], nuclear magnetic resonance spectroscopy [24, 29], seismic analysis [34, 16], system identification [21, 31], time series forecasting [22, 32, 17], and many other real-world applications. Mathematically, a complex-valued Hankel matrix is structured as:
$\bm{X}=\begin{bmatrix}x_{1}&x_{2}&\cdots&x_{n_{2}}\\ x_{2}&x_{3}&\cdots&x_{n_{2}+1}\\ \vdots&\vdots&\cdots&\vdots\\ x_{n_{1}}&x_{n_{1}+1}&\cdots&x_{n}\end{bmatrix}\in\mathbb{C}^{n_{1}\times n_{2}}.$
Since each skew-diagonal of a Hankel matrix has the same values, it contains only $n=n_{1}+n_{2}-1$ distinct values in total. In practice, keeping track of a memory-efficient reweighed vector
$\bm{z}=[x_{1},\sqrt{2}x_{2},\cdots,\sqrt{\varsigma_{i}}x_{i},\cdots,\sqrt{2}x_{n-1},x_{n}]^{\top}\in\mathbb{C}^{n}$
is sufficient to represent a large-scale Hankel matrix, where $\varsigma_{i}$ is the number of entries on the $i$-th antidiagonal of an $n_{1}\times n_{2}$ matrix. Note that $\|\bm{z}\|_{2}=\|\bm{X}\|_{\mathrm{F}}$ which is convenient for formulating the Hankel recovery problem in its equivalent vector form. We also introduce a reweighed Hankel mapping $\mathcal{G}:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n_{1}\times n_{2}}$ and a reweighting operator $\mathcal{W}:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$ such that
$\mathcal{G}\bm{z}=\bm{X}\quad\textnormal{ and }\quad\mathcal{W}\bm{z}=[x_{1},x_{2},\cdots,x_{i},\cdots,x_{n-1},x_{n}]^{\top}.$
Note $\mathcal{G}^{*}\mathcal{G}=\mathcal{I}$ where $(\cdot)^{*}$ denotes the adjoint operator and $\mathcal{I}$ is the identity operator.
In the applications of the low-rank Hankel matrix, we face three major challenges: (i) missing data, (ii) sparse outliers, and (iii) ill condition. To address the first two challenges, we can formulate the Hankel matrix recovery problem:
(1) $\begin{split}\operatorname*{\mathrm{minimize}}_{{\bm{X}},{\bm{O}}\in\mathbb{C}^{n_{1}\times n_{2}}}&~{}\frac{1}{2p}\left\langle\mathcal{P}_{\bm{H}_{\Omega}}{\bm{Y}}-\mathcal{P}_{\bm{H}_{\Omega}}({\bm{X}}+{\bm{O}}),{\bm{Y}}-({\bm{X}}+{\bm{O}})\right\rangle\cr\operatorname*{\mathrm{subject~{}to}}&~{}{\bm{X}}\textnormal{ is rank-$r$ Hankel matrix and }{\bm{O}}\textnormal{ is sparse,}\end{split}$
where $p$ is the observation rate, ${\bm{O}}$ is the sparse outlier matrix, $\mathcal{P}_{\bm{H}_{\Omega}}:\mathbb{C}^{n_{1}\times n_{2}}\rightarrow\mathbb{C}^{n_{1}\times n_{2}}$ is the matrix observation operator, and $\mathcal{P}_{\bm{H}_{\Omega}}{\bm{Y}}$ is the partially observed data matrix with sparse corruptions. However, as discussed in the literature [4], randomly positioned, non-structured missing entries and outliers can be easily recovered through the redundant skew-diagonal in the Hankel matrix. One can also Hankelize $\mathcal{P}_{\bm{H}_{\Omega}}{\bm{Y}}$ by averaging its skew-diagonal elements as ${\bm{X}}$ is Hankel structured. Thus, we are only interested in the more challenging recovery problem with Hankel structured missing entries and outliers, i.e., some skew-diagonals are entirely missed or corrupted. Under this setting, all parties of (1) are Hankel structured including the observation pattern $\bm{H}_{\Omega}$, thus we can rewrite the Hankel recovery problem in the equivalent vector form:
(2) $\begin{split}\operatorname*{\mathrm{minimize}}_{\bm{z},\bm{s}\in\mathbb{C}^{n}}\ &~{}\frac{1}{2p}\left\langle\Pi_{\Omega}\bm{f}-\Pi_{\Omega}(\bm{z}+\bm{s}),\bm{f}-(\bm{z}+\bm{s})\right\rangle\cr\operatorname*{\mathrm{subject~{}to}}&~{}\mathrm{rank}\left(\mathcal{G}\bm{z}\right)=r,\quad\Pi_{\Omega}\bm{s}\textnormal{ is }\alpha\textnormal{-sparse},\end{split}$
where $\Pi_{\Omega}:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$ is the observation operator on vector space that corresponds to the Hankel structured sampling pattern, $\bm{f}=\mathcal{G}^{*}{\bm{Y}}$ and $\bm{s}=\mathcal{G}^{*}{\bm{O}}$ are observed data and outlier matrices in their equivalent vector form, and $\alpha$ is a sparsity parameter defined later in Section 1.1. Notice that only the observed outliers will interfere with the recovery, thus we put the sparsity constraint on $\Pi_{\Omega}\bm{s}$ instead of the entire $\bm{s}$.
Since the applications of Hankel recovery are often time-sensitive, developing efficient solvers is a crucial task. Although many fast non-convex algorithms have been proposed in recent years, they often struggle with ill-conditioned problems (detailed in Section 1.3) which is one of the aforementioned major challenges in applications. This paper aims to study a provable non-convex algorithm for solving (2) that keeps high efficiency even if the condition number of the underlying Hankel matrix is large.
1.1 Assumptions
Denote the vector-form groundtruth as $\bm{z}_{\star}$ and $\bm{s}_{\star}$ for low-rank Hankel and sparse outlier respectively. To address the Hankel recovery problem (2), we introduce the following standard assumptions on data sampling, Hankel matrix incoherence, and outlier sparsity.
{assumption}
[Uniform sampling with replacement] On the vector space $\mathbb{C}^{n}$, the observation index set $\Omega=\{a_{k}\}_{k=1}^{m}$ is sampled uniformly with replacement from $[n]$. The observation rate is defined as $p:=|\Omega|/n=m/n$.
In this context, the sampling operator is defined as $\Pi_{\Omega}\bm{v}=\sum_{i\in\Omega}v_{i}\bm{e}_{i}$ and allows repeated samples, thus $\left\langle\Pi_{\Omega}\bm{v},\bm{v}\right\rangle\neq\|\Pi_{\Omega}\bm{v}\|_{2}^{2}$.
{assumption}
[$\mu$-incoherence] Let ${\bm{U}}_{\star}\bm{\Sigma}_{\star}{\bm{V}}_{\star}^{*}$ be the compact singular value decomposition (SVD) of the underlying rank-$r$ Hankel matrix $\mathcal{G}\bm{z}_{\star}\in\mathbb{C}^{n_{1}\times n_{2}}$. Then $\mathcal{G}\bm{z}_{\star}$ is said $\mu$-incoherent if there exists a positive constant $\mu$ such that
$\|{\bm{U}}_{\star}\|_{2,\infty}\leq\sqrt{\mu c_{s}r/n}\quad\textnormal{ and }\quad\|{\bm{V}}_{\star}\|_{2,\infty}\leq\sqrt{\mu c_{s}r/n},$
where $c_{s}=\max\{{n}/{n_{1}},{n}/{n_{2}}\}$ measures how square the Hankel matrix is.
{assumption}
[$\alpha p$-sparsity] Given a sampling model (i.e., Section 1.1), the observed outliers are at most $\alpha p$-sparse. That is,
$\|\Pi_{\Omega}\bm{s}_{\star}\|_{0}\leq\alpha|\Omega|=\alpha pn.$
Sections 1.1, 1.1 and 1.1 have been widely used in matrix completion [14, 30, 35, 9], robust principal component analysis [13, 28, 3, 10], and low-rank Hankel recovery literature [18, 11, 12, 2]. Together, they ensure that (2) is a well-posed problem [15]. Note that Section 1.1 states only a $\alpha$ proportion of observed data are corrupted, without specifying the randomness of the corruption pattern. This condition holds with high probability as long as $\bm{s}_{\star}$ is $\frac{\alpha}{2}$-sparse [7]. Moreover, the sparsity of $\bm{s}_{\star}$ (resp. $\Pi_{\Omega}\bm{s}_{\star}$) implies the sparsity of $\mathcal{G}\bm{s}_{\star}$ (resp. $\mathcal{G}\Pi_{\Omega}\bm{s}_{\star}$) in each of its rows and columns, provided $n_{1}\approx n_{2}$.
1.2 Notation
[…]
1.3 Related work and contributions
The low-rank Hankel recovery problem has been extensively studied in the past decade. A naive approach is to directly apply some robust matrix recovery algorithms [19, 36, 20, 6, 7, 23, 5, 8] on the Hankel matrix which suffers from high computational complexities and inferior recoverability since the Hankel structure is ignored. Lately, many dedicated Hankel recovery algorithms have been proposed. Robust-EMaC [18] proposes a convex relaxation for the non-convex problem (1) and solves it with semidefinite programming which is still computationally too expensive. [11, 12, 33] propose three algorithms with linear convergence that solve the non-convex Hankel completion problem in $\mathcal{O}(r^{2}n+rn\log n)$ flops; however, none of them handles outliers and [33] also lacks theoretical guarantee. SAP [37] proposes to effectively solve (2) via alternating projections. The recovery guarantee with linear convergence is established with $m\gtrsim\mathcal{O}(c_{s}^{2}\mu^{2}r^{3}\log^{2}n)$ and $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r))$. Note that iterative resample is required for SAP in theory. The computational complexity of SAP is $\mathcal{O}(r^{2}n\log n)$ flops and the accelerated version ASAP [2] improves it to $\mathcal{O}(r^{2}n+rn\log n)$ flops which however only studies the fully observed cases. HSGD [4] solves (2) via projected gradient descent in $\mathcal{O}(r^{2}n+rn\log n)$ flops. Although HSGD has a guaranteed linear convergence, it requires $\mathcal{O}(\kappa\log(1/\varepsilon))$ iterations to find a $\varepsilon$-solution, which is problematic for ill-conditioned problems. The theorem is established with $m\gtrsim\mathcal{O}(\max\{c_{s}^{2}\mu^{2}r^{2}\log n,c_{s}\mu\kappa^{3}r^{2}\log n\})$ and $\alpha\lesssim\mathcal{O}(1/(\max\{c_{s}\mu\kappa^{3/2}r^{3/2},c_{s}\mu r\kappa^{2}\}))$.
In this paper, our main contributions are three-fold:
  1. (i)
    We propose a highly efficient algorithm for ill-conditioned robust Hankel recovery problems, coined Hankel Structured Newton-Like Descent (HSNLD). To find a $\varepsilon$-solution, its computational cost is $\mathcal{O}((r^{2}n+rn\log n)\log(1/\varepsilon))$ flops which beats all existing state-of-the-arts. Especially when compared with HSGD, our convergence rate is independent of the condition number.
  2. (ii)
    The recovery guarantee, with linear convergence, has been established for HSNLD under some mild conditions. The sample complexity and outlier tolerance are $m\gtrsim\mathcal{O}(c_{s}^{2}\mu^{2}r^{2}\kappa^{4}\log n)$ and $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$ respectively.
  3. (iii)
    We verify the empirical advantages of HSNLD with synthetic data and real NMR signal. HSNLD achieves the best recoverability and computational efficiency in all ill-conditioned test problems. It gains more advantages when the condition numbers become even worse.
2 Proposed method
Hankel Structured Gradient Descent (HSGD) [4] is a state-of-the-art non-convex algorithm for Hankel matrix recovery problems. It has shown superior performance for well-conditioned Hankel matrices, both theoretically and empirically. However, when the underlying Hankel matrix is ill-conditioned, the convergence of HSGD is significantly slower in theory. In particular, HSGD requires $\mathcal{O}(\kappa\log\varepsilon^{-1})$ iterations to find a $\varepsilon$-solution. This phenomenon is empirically verified in Fig. 1, where we observe that HSGD suffers from the larger condition number of the underlying Hankel matrix. To overcome this major weakness of HSGD, we propose a novel non-convex algorithm inspired by recent developments on Newton-like Gradient Descent algorithms [33]. The proposed algorithm, dubbed Hankel Structured Newton-like Descent (HSNLD), is highly efficient even for ill-conditioned problems.
Figure 1: The convergence performance of HSGD [4] with different condition numbers $\kappa$.
To recover the low-rank Hankel matrix from missing data and sparse outliers, we aim at a similar objective function as used in [4]:
(3) $\begin{split}\ell\left({\bm{L}},{\bm{R}};\bm{s}\right):=&~{}\frac{1}{2p}\left\langle\Pi_{\Omega}\left(\mathcal{G}^{*}\left({\bm{L}}{\bm{R}}^{*}\right)+\bm{s}-\bm{f}\right),\mathcal{G}^{*}\left({\bm{L}}{\bm{R}}^{*}\right)+\bm{s}-\bm{f}\right\rangle\cr&~{}+\frac{1}{2}\left\lVert\left(\mathcal{I}-\mathcal{G}\mathcal{G}^{*}\right)\left({\bm{L}}{\bm{R}}^{*}\right)\right\lVert_{\mathrm{F}}^{2},\end{split}$
where $\bm{z}=\mathcal{G}^{*}{\bm{X}}\in\mathbb{C}^{n}$ and we parameterize the low-rank Hankel matrix ${\bm{X}}\in\mathbb{C}^{n_{1}\times n_{2}}$ as a product of two tall matrices ${\bm{L}}\in\mathbb{C}^{n_{1}\times r}$ and ${\bm{R}}\in\mathbb{C}^{n_{2}\times r}$, thus the low-rank constraint is implicitly encoded. Moreover, a regularization term $\left\lVert\left(\mathcal{I}-\mathcal{G}\mathcal{G}^{*}\right)\left({\bm{L}}{\bm{R}}^{*}\right)\right\lVert_{\mathrm{F}}^{2}$ is added to enforce the Hankel structure of ${\bm{L}}{\bm{R}}^{*}$. Targeting (3), the proposed HSNLD (summarized in Algorithm 1) will be discussed step by step in the following paragraphs. For the sake of presentation, the iterative steps will be discussed first, and then initialization.
Remark 2.1.
Compared to the objective posted in [4], our objective (3) omits the balance regularization term $\|{\bm{L}}^{*}{\bm{L}}-{\bm{R}}^{*}{\bm{R}}\|_{\mathrm{F}}^{2}$. The balance regularization is not empirically needed for the original HSGD algorithm but is required for its convergence analysis. In this paper, we drop the balance regularization for the proposed HSNLD with an improved convergence guarantee.
Algorithm 1 Hankel Structured Newton-Like Descent (HSNLD)
1:Input: $\Pi_{\Omega}\bm{f}$: partial observation on the corrupted Hankel matrix in reweighted vector form; $r$: the rank of underlying Hankel matrix; $p$: observation rate; $\alpha$: outlier density; $\{\gamma_{k}\}$: parameters for sparsification operator; $K$: number of iterations;
2:Initialization: Let ${\bm{L}}_{0},{\bm{R}}_{0}$ be produced by (7).
3:for  $k=0,1,\ldots,K-1$  do
4:     $\bm{z}_{k+1}=\mathcal{G}^{*}({\bm{L}}_{k}{\bm{R}}_{k}^{*})$
5:     $\bm{s}_{k+1}=\Gamma_{\gamma_{k}\alpha p}(\Pi_{\Omega}(\bm{f}-\bm{z}_{k+1}))$
6:     $\begin{bmatrix}{\bm{L}}_{k+1}\\ {\bm{R}}_{k+1}\end{bmatrix}=\mathcal{P}_{\mathcal{C}}\begin{bmatrix}{\bm{L}}_{k}-\eta\nabla_{{\bm{L}}}\ell({\bm{L}}_{k},{\bm{R}}_{k};\bm{s}_{k+1})\left({\bm{R}}^{*}_{k}{\bm{R}}_{k}\right)^{-1}\\ {\bm{R}}_{k}-\eta\nabla_{{\bm{R}}}\ell({\bm{L}}_{k},{\bm{R}}_{k};\bm{s}_{k+1})\left({\bm{L}}^{*}_{k}{\bm{L}}_{k}\right)^{-1}\end{bmatrix}$
7:end for
8:Output: $\bm{z}_{\mathrm{output}}=\mathcal{G}^{*}({\bm{L}}_{K}{\bm{R}}_{K}^{*})$ : recovered Hankel matrix in vector form.
Iterative updates on $\bm{s}$. For the sake of computational efficiency, one should avoid forming the full-size Hankel matrix and process the calculation in the corresponding vector form. Hence, we first update $\bm{z}$ with the current low-rank components, i.e.,
(4) $\bm{z}_{k+1}=\mathcal{G}^{*}({\bm{L}}_{k}{\bm{R}}_{k}^{*}).$
By the definition of $\mathcal{G}^{*}$, the $t$-th entry of $\bm{z}$ can be written as
$[\mathcal{G}^{*}({\bm{L}}{\bm{R}}^{*})]_{t}=\sum_{j=1}^{r}[\mathcal{G}^{*}({\bm{L}}_{:,j}{\bm{R}}_{:,j}^{*})]_{t}=\sum_{j=1}^{r}\frac{1}{\sqrt{\varsigma_{t}}}\sum_{i_{1}+i_{2}=t+1}L_{i_{1},j}\overline{R}_{i_{2},j}.$
Thus, (4) can be obtained via $r$ fast convolutions, with a low computational cost in $\mathcal{O}(rn\log n)$ flops.
To detect the outliers, we compute the residue $\Pi_{\Omega}(\bm{f}-\bm{z}_{k+1})$ over the observation and project it onto the space of sparse vectors via a sparsification operator. That is,
(5) $\bm{s}_{k+1}=\Gamma_{\gamma_{k}\alpha p}\Pi_{\Omega}\big{(}\bm{f}-\bm{z}_{k+1}\big{)},$
where $\Gamma_{\gamma_{k}\alpha p}:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$ is the vector sparsification operator that keeps the $\gamma_{k}\alpha pn$ largest-in-magnitude entries unchanged and vanishes the rest to 00. Therein, $\gamma_{k}>1$ is a parameter that tolerates slight overestimation on the outlier sparsity as the true $\alpha$ is usually unavailable to the user, and it also provides redundant robustness in outlier detection. We observe that taking $\gamma_{k}\rightarrow 1$ as $k\rightarrow\infty$ provides balanced performance in robustness and efficiency; however, $\gamma_{k}$ should always be larger than $1$. The computational cost of (5) is $\mathcal{O}(n\log n)$ flops or better.
Iterative updates on ${\bm{L}}$ and ${\bm{R}}$. The state-of-the-art HSGD [4] updates the low-rank Hankel components via iterative gradient descent. As discussed, the convergence of HSGD suffers from ill-conditioned problems. Inspired by the recent development in [33], we utilize an additional Newton-like preconditioner in the iterative updates of ${\bm{L}}$ and ${\bm{R}}$ to properly address the ill-conditioned Hankel matrices. We have
(6) $\begin{bmatrix}{\bm{L}}_{k+1}\\ {\bm{R}}_{k+1}\end{bmatrix}=\mathcal{P}_{\mathcal{C}}\begin{bmatrix}{\bm{L}}_{k}-\eta\nabla_{{\bm{L}}}\ell({\bm{L}}_{k},{\bm{R}}_{k};\bm{s}_{k+1})\left({\bm{R}}^{*}_{k}{\bm{R}}_{k}\right)^{-1}\\ {\bm{R}}_{k}-\eta\nabla_{{\bm{R}}}\ell({\bm{L}}_{k},{\bm{R}}_{k};\bm{s}_{k+1})\left({\bm{L}}^{*}_{k}{\bm{L}}_{k}\right)^{-1}\end{bmatrix}$
where $\eta>0$ is the step size and $\mathcal{P}_{\mathcal{C}}$ is a projection operator that enforces the weak incoherence condition, i.e., $\max\{\|{\bm{L}}{\bm{R}}^{*}\|_{2,\infty},\|{\bm{R}}{\bm{L}}^{*}\|_{2,\infty}\}\leq C$, on the low-rank Hankel matrix. In particular, modified from [33] we define
$\mathcal{P}_{\mathcal{C}}\begin{bmatrix}\widetilde{{\bm{L}}}\\ \widetilde{{\bm{R}}}\end{bmatrix}=\begin{bmatrix}{\bm{L}}\\ {\bm{R}}\end{bmatrix},~{}\text{where}~{}{\bm{L}}_{i,:}$ $=\min\left(1,{C}/{\big{\|}\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{\frac{1}{2}}\big{\|}_{2}}\right)\widetilde{{\bm{L}}}_{i,:},\quad\forall i\in[n_{1}],$
${\bm{R}}_{j,:}$ $=\min\left(1,{C}/{\big{\|}\widetilde{{\bm{R}}}_{j,:}(\widetilde{{\bm{L}}}^{*}\widetilde{{\bm{L}}})^{\frac{1}{2}}\big{\|}_{2}}\right)\widetilde{{\bm{R}}}_{j,:},\quad\forall j\in[n_{2}],$
for some numerical constant $C>0$. The choice of $C$ will be specified in Section 3 and the locally non-expansiveness of $\mathcal{P}_{\mathcal{C}}$ is verified in Lemma 5.12.
To efficiently update ${\bm{L}}$ and ${\bm{R}}$, we notice that
$\quad~{}\nabla_{\bm{L}}\ell({\bm{L}},{\bm{R}};\bm{s})({\bm{R}}^{*}{\bm{R}})^{-1}$
$=\left[\mathcal{G}\big{(}p^{-1}\Pi_{\Omega}\left(\mathcal{G}^{*}\left({\bm{L}}{\bm{R}}^{*}\right)+\bm{s}-\bm{f}\right)\big{)}{\bm{R}}+\left({\bm{I}}-\mathcal{G}\mathcal{G}^{*}\right)({\bm{L}}{\bm{R}}^{*}){\bm{R}}\right]({\bm{R}}^{*}{\bm{R}})^{-1}$
$=\mathcal{G}\big{(}p^{-1}\Pi_{\Omega}\left(\mathcal{G}^{*}\left({\bm{L}}{\bm{R}}^{*}\right)+\bm{s}-\bm{f}\right)-\mathcal{G}^{*}({\bm{L}}{\bm{R}}^{*})\big{)}{\bm{R}}({\bm{R}}^{*}{\bm{R}})^{-1}+{\bm{L}}.$
One can compute $\bm{z}:=\mathcal{G}^{*}({\bm{L}}{\bm{R}}^{*})$ via $r$ fast convolutions which cost $\mathcal{O}(rn\log n)$ flops. Computing $p^{-1}\Pi_{\Omega}\left(\bm{z}+\bm{s}-\bm{f}\right)-\bm{z}$ costs $\mathcal{O}(n)$ flops or better. Since $\mathcal{G}(\cdot){\bm{R}}$ can be viewed as $r$ convolutions, it costs $\mathcal{O}(rn\log n)$ flops. Computing ${\bm{R}}({\bm{R}}^{*}{\bm{R}})^{-1}$ costs $\mathcal{O}(r^{2}n)$ flops, so does the projection $\mathcal{P}_{\mathcal{C}}$. Thus, the total computational cost of updating ${\bm{L}}$ (and similarly for ${\bm{R}}$) is $\mathcal{O}(rn\log n+r^{2}n)$
Initialization. A tight initialization is important to the success of HSNLD. We adopt the modified Hankel spectral initialization from [4]:
(7) $\begin{split}\bm{s}_{0}&=\mathcal{W}^{-1}\Gamma_{\alpha p}\left(\mathcal{W}\Pi_{\Omega}\bm{f}\right),\cr{\bm{U}}_{0}\bm{\Sigma}_{0}{\bm{V}}_{0}^{*}&=\mathrm{SVD}_{r}(p^{-1}\mathcal{G}(\Pi_{\Omega}\bm{f}-\bm{s}_{0})),\cr\begin{bmatrix}{\bm{L}}_{0}\\ {\bm{R}}_{0}\end{bmatrix}&=\mathcal{P}_{\mathcal{C}}\begin{bmatrix}{\bm{U}}_{0}\bm{\Sigma}_{0}^{1/2}\\ {\bm{V}}_{0}\bm{\Sigma}_{0}^{1/2}\end{bmatrix}.\end{split}$
This initialization has shown promising performance and a low computational complexity of $\mathcal{O}(rn\log n)$. Note that the truncated SVD of a Hankel matrix has a low cost since all matrix-vector multiplications therein can be viewed as convolutions.
Therefore, the overall computational complexity of the proposed algorithm is $\mathcal{O}(rn\log n+r^{2}n)$ which is tied to the current state-of-the-art HSGD [4]. In addition, we emphasize that HSNLD never needs to form the entire Hankel matrix and that tracking the $n$ distinct elements in the corresponding vector form is sufficient. Thus, the proposed HSNLD is computationally and memory efficient.
3 Theoretical results
We present the theoretical results of the proposed HSNLD in this section. We will first define some convenient notations and metrics for the presentation. Denote $\bm{L}_{\star}:={\bm{U}}_{\star}\bm{\Sigma}_{\star}^{1/2}$ and $\bm{R}_{\star}:={\bm{V}}_{\star}\bm{\Sigma}_{\star}^{1/2}$ where ${\bm{U}}_{\star}\bm{\Sigma}_{\star}{\bm{V}}_{\star}^{*}=\mathcal{G}\bm{z}_{\star}$ is the compact SVD of the underlying Hankel matrix. Define the distance between $({\bm{L}}_{k},{\bm{R}}_{k})$ and $(\bm{L}_{\star},\bm{R}_{\star})$ to be
(8) $d_{k}^{2}:=$ $\mathrm{dist}^{2}({\bm{L}}_{k},{\bm{R}}_{k};\bm{L}_{\star},\bm{R}_{\star})$
(9) ::=:= $\inf_{{\bm{Q}}\in\mathrm{GL}(r,\mathbb{C})}\big{\|}\left({\bm{L}}_{k}{\bm{Q}}-\bm{L}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\big{\|}_{\mathrm{F}}^{2}+\big{\|}\left({\bm{R}}_{k}{\bm{Q}}^{-*}-\bm{R}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\big{\|}_{\mathrm{F}}^{2}.$
Therein, $\mathrm{GL}(r,\mathbb{C})$ denotes the general linear group of degree $r$ over $\mathbb{C}$ and the invertible matrix ${\bm{Q}}$ is the best alignment between $({\bm{L}}_{k},{\bm{R}}_{k})$ and $(\bm{L}_{\star},\bm{R}_{\star})$. [33, Lemma 22] has shown the best alignment ${\bm{Q}}$ always exists if $d_{k}$ is sufficiently small. In that case, the infimum in (8) can be replaced by the minimum, and the standard recovery error $\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right\lVert_{\mathrm{F}}$ can be bounded by $d_{k}$ as stated in the following proposition.
Proposition 3.1.
For any $\varepsilon_{0}\in(0,1)$, if provided $d_{k}<\varepsilon_{0}\sigma^{\star}_{r}$, then it holds that
$\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right\lVert_{\mathrm{F}}\leq\sqrt{2}(1+\varepsilon_{0}/2)d_{k}.$
Proof 3.2.
The proof of this proposition is deferred to Section 5.2.
The following Theorem 3.3 guarantees that our initialization (7) will start HSNLD sufficiently close to the ground truth, under some mild conditions. Thus, it is sufficient to bound $d_{k}$ for the recovery guarantee. In addition, our initial guess also satisfies the weak incoherence condition, which will be used and inherited iteratively in the convergence analysis. For ease of presentation, we denote the basin of attraction as:
$\mathcal{E}(\delta,C):=\{({\bm{L}},{\bm{R}})~{}|~{}\mathrm{dist}({\bm{L}},{\bm{R}};\bm{L}_{\star},\bm{R}_{\star})\leq\delta~{}~{}\textnormal{and}~{}~{}\max\{\left\lVert{\bm{L}}{\bm{R}}^{*}\right\lVert_{2,\infty},\left\lVert{\bm{R}}{\bm{L}}^{*}\right\lVert_{2,\infty}\}\leq C\}.$
Theorem 3.3.
For any $\varepsilon_{0}\in(0,1)$, let $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1+\varepsilon_{0}$, and suppose Sections 1.1, 1.1 and 1.1 hold with $m\geq c_{0}\varepsilon_{0}^{-2}c_{s}\mu\kappa^{2}r^{2}\log n$ and $\alpha\leq\varepsilon_{0}/(50c_{s}\mu r^{1.5}\kappa)$, where $c_{0}$ is some universal constant. Then, the initialization step (7) satisfies
$({\bm{L}}_{0},{\bm{R}}_{0})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},C),$
with probability at least $1-2n^{-2}$. In the case of full observation, probability rises to $1$.
Proof 3.4.
The proof of this theorem is deferred to Section 5.3.
3.1 Full observation
We start the analysis with the simpler full observation case, i.e., $\Omega=[n]$. Theorem 3.5 presents the local linear convergence of HSNLD.
Theorem 3.5.
Suppose Sections 1.1 and 1.1 hold with $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$. Choose parameters $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1.01$, $\gamma_{k}\in[1+1/b_{0},2]$ with some fixed $b_{0}\geq 1$, and $\eta\in(0,1]$. Under full observation, for any $({\bm{L}}_{k},{\bm{R}}_{k})\in\mathcal{E}(0.01\sigma^{\star}_{r},C)$, it holds
$d_{k+1}\leq(1-0.6\eta)d_{k}\qquad\textnormal{and}\qquad({\bm{L}}_{k+1},{\bm{R}}_{k+1})\in\mathcal{E}(0.01\sigma^{\star}_{r},C).$
Proof 3.6.
The proof of this theorem is deferred to Section 5.4.1.
Combining Propositions 3.1, 3.3 and 3.5, we have the recovery guarantee, with a linear convergence, for the proposed HSNLD under full observation.
Corollary 3.7 (Recovery guarantee with full observation).
Suppose Sections 1.1 and 1.1 hold with $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$. Choose parameters $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1.01$, $\gamma_{k}\in[1+1/b_{0},2]$ with some fixed $b_{0}\geq 1$, and $\eta\in(0,1]$. Then, under full observation, in $K=\mathcal{O}(\log(1/\varepsilon))$ iterations, the outputs of Algorithm 1 satisfy
$\left\lVert{\bm{L}}_{K}{\bm{R}}_{K}^{*}-\mathcal{G}\bm{z}_{\star}\right\lVert_{\mathrm{F}}\leq 0.02(1-0.6\eta)^{K}\sigma^{\star}_{r}\leq\varepsilon.$
3.2 Partial observation
Now we will process the partial observation cases. One major challenge is the dependency among the entries on the antidiagonals of the Hankel matrix, which causes difficulty in uniformly bounding the subsampling error (see Lemma 5.26). Now, we present the local convergence under partial observation.
Theorem 3.8.
Suppose Sections 1.1, 1.1 and 1.1 hold with $m\gtrsim\mathcal{O}(c_{s}^{2}\mu^{2}r^{2}\kappa^{4}\log n)$ and $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$. Choose the parameters $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1.01$, $\gamma_{k}\in[1+1/b_{0},2]$ with some fixed $b_{0}\geq 1$, and $\eta\in(0,0.6]$. For any fixed $({\bm{L}}_{k},{\bm{R}}_{k})\in\mathcal{E}(0.01\sigma^{\star}_{r},C)$, with probability at least $1-8n^{-2}$ the following results hold
$d_{k+1}\leq(1-0.6\eta)d_{k}\qquad\textnormal{and}\qquad({\bm{L}}_{k+1},{\bm{R}}_{k+1})\in\mathcal{E}(0.01\sigma^{\star}_{r},C).$
Proof 3.9.
The proof of this theorem is deferred to Section 5.4.2.
Combining Propositions 3.1, 3.3 and 3.8, we have the recovery guarantee, with a linear convergence, for the proposed HSNLD under partial observation.
Corollary 3.10 (Recovery guarantee with partial observation).
Suppose Sections 1.1, 1.1 and 1.1 hold with $m\gtrsim\mathcal{O}(c_{s}^{2}\mu^{2}r^{2}\kappa^{4}\log n)$ and $\alpha\lesssim\mathcal{O}(1/(c_{s}\mu r\kappa^{2}))$. Let the observation set $\Omega$ be resampled independently every iteration. Choose parameters $C=c\sqrt{c_{s}\mu r/n}\sigma^{\star}_{1}$ with $c\geq 1.01$, $\gamma_{k}\in[1+1/b_{0},2]$ with some fixed $b_{0}\geq 1$, and $\eta\in(0,0.6]$. Then, in $K=\mathcal{O}(\log(1/\varepsilon))$ iterations, the outputs of Algorithm 1 satisfy
$\left\lVert{\bm{L}}_{K}{\bm{R}}_{K}^{*}-\mathcal{G}\bm{z}_{\star}\right\lVert_{\mathrm{F}}\leq 0.02(1-0.6\eta)^{K}\sigma^{\star}_{r}\leq\varepsilon$
with probability at least $1-\mathcal{O}(n^{-1})$, provided $\log(1/\varepsilon)\ll n$.
Remark 3.11.
Compared to the state-of-the-art HSGD which runs $\mathcal{O}(\kappa\log(1/\varepsilon))$ iterations for a $\varepsilon$-optimizer, the convergence rate of HSNLD is independent of the condition number $\kappa$. This makes HSNLD a powerful tool for ill-conditioned Hankel applications, e.g., spectrally sparse signal recovery with inseparable frequencies [27].
4 Numerical experiments
In this section, we will empirically demonstrate the performance of the proposed HSNLD with synthetic and real datasets. We compare our algorithm with HSGD [4], SAP [37], and RobustEMaC [18]. Parameters are set according to the original paper and further hand-tuned for best performance when applicable. Particularly, both HSNLD and HSGD use the same outlier sparsification parameters: $\gamma_{k}=1.05+0.45\times 0.95^{k}$ as suggested in [4]. Moreover, we estimate the incoherence parameter $\mu$ via one-step Cadzow method [1] for all tested algorithms. The resampling is not implemented for HSNLD as it is just a proof technique and is not needed in practice. All results were conducted on Matlab with Intel i9-12950HX and 64GB RAM. The code is available online at https://github.com/caesarcai/HSNLD.
4.1 Synthetic dataset
We generate an ill-conditioned Hankal matrix in its vector form $\bm{z}\in\mathbb{C}^{n}$ with $z_{t}=\sqrt{\varsigma_{t}}\sum^{r}_{j=1}a_{j}e^{2\pi\imath f_{j}t},$ where $\imath$ denotes the imaginary unit, $\{f_{j}\}$ is a randomly selected set of normalized off-grid frequencies, and $\{a_{j}\}$ is a set of $r$ evenly spaced numbers over $[1/\kappa,1]$. Thus, $\mathcal{G}\bm{z}\in\mathbb{C}^{n_{1}\times n_{2}}$ is exactly rank-$r$ with condition number $\kappa$. Then, $m$ observations are sampled uniformly without replacement from $\bm{z}$, and therein $\alpha m$ entries are randomly chosen for corruption. We add complex outliers to the corrupted entries whose real and imaginary parts are drawn uniformly over the intervals $[-10\mathbb{E}(|\mathrm{Re}(z_{t})|),10\mathbb{E}(|\mathrm{Re}(z_{t})|)]$ and $[-10\mathbb{E}(|\mathrm{Im}(z_{t})|),10\mathbb{E}(|\mathrm{Im}(z_{t})|)]$.
Phase transition. We demonstrate the recoverability and robustness of HSNLD via empirical phase transition under various settings. When an algorithm stops at $\|\mathcal{G}\Pi_{\Omega}\bm{z}_{k}+\mathcal{G}\Pi_{\Omega}\bm{s}_{k}-\mathcal{G}\Pi_{\Omega}\bm{f}\|_{\mathrm{F}}/\|\mathcal{G}\Pi_{\Omega}\bm{f}\|_{\mathrm{F}}\leq 10^{-5}$, we considered it a successful recovery if $\|\mathcal{G}\bm{z}_{k}-\mathcal{G}\bm{z}_{\star}\|_{\mathrm{F}}/\|\mathcal{G}\bm{z}_{\star}\|_{\mathrm{F}}\leq 10^{-3}$. For every problem setting, each algorithm runs for $20$ trials and reports the result as a pixel on the phase transition, where a white pixel means all $20$ trials were recovered and a black pixel means all $20$ trials were failed. The phase transition results are reported in Figs. 2, 4 and 3 where one can see HSNLD delivers the best recoverability and robustness for ill-conditioned problems.
Figure 2: Phase transition: Number of samples vs. rate of outliers. Problem dimension $n=125$, condition number $\kappa=10$, and rank $r=10$ for every problem.
Figure 3: Phase transition: Rank vs. rate of outliers. Problem dimension $n=125$, condition number $\kappa=10$, and $m=50$ entries are sampled in every problem.
Figure 4: Phase transition: Number of samples vs. rank. Problem dimension $n=125$, condition number $\kappa=10$, and outlier rate $\alpha=10\%$ for every problem.
Convergence performance. In Fig. 5, we plot the relative error v.s. runtime for the tested algorithms with different condition numbers. All problems have dimension $n=2^{16}-1$ and the results are averaged over 5 trials. When the condition number is perfect, i.e., $\kappa=1$, the HSNLD has tied convergence performance with HSGD. When the condition number grows, HSGD’s performance drops dramatically while HSNLD takes no impact—this matches our theoretical results. Generally speaking, SAP’s convergence performance is not impacted by the condition number either; however, the truncated SVD it applies every iteration may fail to function with vast condition number. RobustEMaC is not tested here because its sublinear convergence is too slow for the comparison.
Figure 5: Convergence performance: Relative error v.s. runtime with condition numbers $\kappa=1,5,20,2000$. Note that the singular values of complex-valued SVD failed to stay real with vast condition number $\kappa=2000$, thus no convergence result can be reported for SAP.
4.2 Nuclear magnetic resonance dataset
We test HSNLD with nuclear magnetic resonance (NMR) signal recovery problem, a classic application of low-rank Hankel recovery. The 1D NMR signal being tested has a length of $32,768$ with an approximate rank of $40$ and condition number $30$. The observation rate is set to be $30\%$ and $10\%$ outliers are added. All algorithms stop at $|\mathrm{err}_{k+1}-\mathrm{err}_{k}|/\mathrm{err}_{k}<10^{-4}$. The recovery results are reported in Figs. 6 and 1. One can see that all three tested algorithms recovered the signal successfully while HSNLD is significantly faster.
Figure 6: NMR recovery results in the form of power spectrum.
Table 1: The number of iterations and runtime for NMR recovery.
Method HSNLD HSGD SAP
Iteration 43 314 40
Runtime 5.22 28.82 79.68
5 Proofs
In this section, we provide the analysis of the claimed theoretical results. All the proofs are under Sections 1.1, 1.1 and 1.1. We start with introducing some additional notations used in the analysis.
For ease of presentation, in addition to Section 1.2, we denote the following slightly abused notation for the proofs. We let $\Omega$ be any subset of $[n]$ that satisfies Section 1.1 with $m:=\lvert\Omega\rvert$ and $p:=m/n$. We denote the tangent space of rank-$r$ matrix manifold at $\mathcal{G}\bm{z}_{\star}$ by
$T:=\left\{\bm{X}~{}|~{}\bm{X}={\bm{U}}_{\star}\bm{C}^{*}+\bm{D}{\bm{V}}_{\star}^{*}\ \text{for any}\ \bm{C}\in\mathbb{C}^{n_{1}\times r},~{}\bm{D}\in\mathbb{C}^{n_{2}\times r}\right\}.$
Let $\Omega_{a}$ denote the index set of the $a$-th antidiagonal of $n_{1}\times n_{2}$ matrix, i.e.,
(10) $\Omega_{a}:=\{(i,j):i+j=a+1,1\leq i\leq n_{1},1\leq j\leq n_{2}\},~{}a=1,2\cdots,n.$
We also introduce the following notations for the ease of presentation: Let ${\bm{Q}}_{k}$ be the best rotation matrix that aligns $({\bm{L}}_{k},{\bm{R}}_{k})$ and $(\bm{L}_{\star},\bm{R}_{\star})$, and according to [33, Lemma 22] we know ${\bm{Q}}_{k}$ always exist if $d_{k}<\sigma^{\star}_{r}$. Denote
$\bm{L}_{\natural}:={\bm{L}}_{k}{\bm{Q}}_{k},~{}\bm{R}_{\natural}:={\bm{R}}_{k}{\bm{Q}}_{k}^{-*},~{}\bm{\Delta}_{{\bm{L}}_{k}}:=\bm{L}_{\natural}-\bm{L}_{\star},~{}\bm{\Delta}_{{\bm{R}}_{k}}:=\bm{R}_{\natural}-\bm{R}_{\star}.$
5.1 Technical lemmas
Lemma 5.1 ([12, Lemma 2][4, Lemma 9]).
It holds with probability at least $1-2n^{-2}$ that
(11) $\left\lVert\left(p^{-1}\mathcal{G}\Pi_{\Omega}-\mathcal{G}\right)\bm{z}_{\star}\right\lVert_{2}\leq\tilde{c}_{0}\sqrt{(\mu c_{s}r\log n)/m}\left\lVert\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$
provided $m\geq\mu c_{s}r\log n$. $\tilde{c}_{0}$ is a universal constant.
Proof 5.2.
The bound of $\left\lVert\left(p^{-1}\mathcal{G}\Pi_{\Omega}-\mathcal{G}\right)\bm{z}_{\star}\right\lVert_{2}$ is provided in [12, Lemma 2] (see [12, Eq. (25)]), and also proved with a slightly different sampling model in [4, Lemma 9].
Lemma 5.3 ([4, Lemma 10]).
There exists a universal constant $\tilde{c}_{1}$ such that if provided $m\geq\varepsilon_{0}^{-2}\tilde{c}_{1}\mu r\log n$, it holds
(12) $\left\lVert p^{-1}\mathcal{P}_{T}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}\mathcal{P}_{T}-\mathcal{P}_{T}\mathcal{G}\mathcal{G}^{*}\mathcal{P}_{T}\right\lVert\leq\varepsilon_{0}$
with probability at least $1-2n^{-2}$.
Proof 5.4.
The result is proved in [4, Lemma 9] under a slightly different sampling model and can be extended to our sampling model, which is omitted here for simplicity.
Lemma 5.5 ([11, Lemma 5]).
For all $\bm{u}\in\mathbb{R}^{n_{1}}$, $\bm{v}\in\mathbb{R}^{n_{2}}$, it holds
(13) $p^{-1}\sum_{a_{k}\in\Omega}\sum_{(i,j)\in\Omega_{a_{k}}}u_{i}v_{j}\leq\left\lVert\bm{u}\right\lVert_{1}\left\lVert\bm{v}\right\lVert_{1}+\sqrt{24np^{-1}\log n}\left\lVert\bm{u}\right\lVert_{2}\left\lVert\bm{v}\right\lVert_{2}$
with probability at least $1-2n^{-2}$, provided $m\geq 3\log n$.
Lemma 5.6.
If $m\geq\log n$, then under event (13) it holds
$~{}p^{-1}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right),\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right\rangle\right\rvert$
$\leq$ $~{}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}+\sqrt{24p^{-1}n\log n}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}},$
for all ${\bm{L}}_{{\bm{A}}}\in\mathbb{C}^{n_{1}\times r}$ and ${\bm{R}}_{{\bm{A}}}\in\mathbb{C}^{n_{2}\times r}$.
Proof 5.7.
For $\Omega=\{a_{k}\}_{k=1}^{m}$, we have
$~{}p^{-1}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right),\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right\rangle\right\rvert$
$=$ $~{}p^{-1}\sum_{k=1}^{m}\bigg{|}\sum_{(i,j)\in\Omega_{a_{k}}}\frac{1}{\sqrt{\varsigma_{a_{k}}}}\left\langle{\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*},\bm{e}_{i}\bm{e}_{j}^{\top}\right\rangle\bigg{|}^{2}$
$\leq$ $~{}p^{-1}\sum_{k=1}^{m}\sum_{(i,j)\in\Omega_{a_{k}}}\frac{1}{\sqrt{\varsigma_{a_{k}}}}\left\lVert{\bm{L}}_{{\bm{A}}}(i,:)\right\lVert_{2}^{2}\left\lVert{\bm{R}}_{{\bm{A}}}(j,:)\right\lVert_{2}^{2}$
$\leq$ $~{}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}+\sqrt{24p^{-1}n\log n}\sqrt{\sum_{i}\left\lVert{\bm{L}}_{{\bm{A}}}(i,:)\right\lVert_{2}^{4}}\sqrt{\sum_{j}\left\lVert{\bm{R}}_{{\bm{A}}}(j,:)\right\lVert_{2}^{4}}$
$\leq$ $~{}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}+\sqrt{24p^{-1}n\log n}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}.$
where the second line follows from Cauchy-Schwarz inequality and the third line follows from Lemma 5.5.
Lemma 5.8.
For all matrices ${\bm{A}},{\bm{B}}\in T$, under event (12), it holds
$\left\lvert\left\langle\mathcal{G}(p^{-1}\Pi_{\Omega}-\mathcal{I})\mathcal{G}^{*}\bm{A},\bm{B}\right\rangle\right\rvert\leq\varepsilon_{0}\left\lVert{\bm{A}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{B}}\right\lVert_{\mathrm{F}},$
and
$\left\lvert\left\langle\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}{\bm{A}},{\bm{A}}\right\rangle\right\rvert\leq p(1+\varepsilon_{0})\left\lVert{\bm{A}}\right\lVert_{\mathrm{F}}^{2}.$
Proof 5.9.
For any matrices ${\bm{A}},{\bm{B}}\in T$, we have $\mathcal{P}_{T}{\bm{A}}={\bm{A}}$ and $\mathcal{P}_{T}{\bm{B}}={\bm{B}}$. Then, under event (12), it holds
$\left\lvert\left\langle\mathcal{G}(p^{-1}\Pi_{\Omega}-\mathcal{I})\mathcal{G}^{*}{\bm{A}},{\bm{B}}\right\rangle\right\rvert=\left\lvert\left\langle\mathcal{P}_{T}\mathcal{G}(p^{-1}\Pi_{\Omega}-\mathcal{I})\mathcal{G}^{*}\mathcal{P}_{T}{\bm{A}},{\bm{B}}\right\rangle\right\rvert\leq\varepsilon_{0}\left\lVert{\bm{A}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{B}}\right\lVert_{\mathrm{F}},$
Thus, together with the inequality $\left\lVert\mathcal{G}^{*}{\bm{A}}\right\lVert_{2}^{2}=\left\langle\mathcal{G}\mathcal{G}^{*}{\bm{A}},{\bm{A}}\right\rangle\leq\left\lVert{\bm{A}}\right\lVert_{\mathrm{F}}^{2}$ we have
$\left\lvert\left\langle\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}{\bm{A}},{\bm{A}}\right\rangle\right\rvert\leq p\varepsilon_{0}\left\lVert{\bm{A}}\right\lVert_{\mathrm{F}}^{2}+p\left\lVert\mathcal{G}^{*}{\bm{A}}\right\lVert_{2}^{2}\leq p(1+\varepsilon_{0})\left\lVert{\bm{A}}\right\lVert_{\mathrm{F}}^{2}.$
Lemma 5.10.
For any $\bm{z}\in\mathbb{C}^{n}$ and $\bm{s}=\mathcal{W}^{-1}\Gamma_{\gamma\alpha p}(\mathcal{W}\Pi_{\Omega}(\bm{f}-\bm{z}))$ with $\gamma\geq 1$, we have
(14) $\big{\|}\mathcal{W}\Pi_{\Omega}\bm{s}_{\star}-\mathcal{W}\bm{s}\big{\|}_{\infty}\leq 2\left\lVert\mathcal{W}\Pi_{\Omega}(\bm{z}_{\star}-\bm{z})\right\lVert_{\infty}.$
Proof 5.11.
Recall $\bm{f}=\bm{z}_{\star}+\bm{s}_{\star}$. By the definition of $\bm{s}$, we have
$\left[\mathcal{W}\Pi_{\Omega}\bm{s}_{\star}-\mathcal{W}\bm{s}\right]_{i}=-\left[\mathcal{W}\Pi_{\Omega}\left(\bm{z}_{\star}-\bm{z}\right)\right]_{i},\quad\text{for}~{}i\in\mathrm{supp}\left(\bm{s}\right).$
Given that $\left\lVert\Pi_{\Omega}\bm{s}_{\star}\right\lVert_{0}\leq\alpha pn=\alpha m$, there can be at most $\alpha m$ elements in $\mathcal{W}\Pi_{\Omega}\left(\bm{f}-\bm{z}\right)$ satisfying $|[\mathcal{W}\Pi_{\Omega}\left(\bm{f}-\bm{z}\right)]_{i}|>\left\lVert\mathcal{W}\Pi_{\Omega}\left(\bm{z}_{\star}-\bm{z}\right)\right\lVert_{\infty}$. Then, for $\gamma\geq 1$ we have
$|[\mathcal{W}\Pi_{\Omega}\left(\bm{f}-\bm{z}\right)]_{i}|\leq\left\lVert\mathcal{W}\Pi_{\Omega}\left(\bm{z}_{\star}-\bm{z}\right)\right\lVert_{\infty},\quad\text{for}~{}i\in\mathrm{supp}(\bm{s}_{\star})\setminus\mathrm{supp}\left(\bm{s}\right).$
Thus, for $i\in\mathrm{supp}(\bm{s}_{\star})\setminus\mathrm{supp}\left(\bm{s}\right)$ we have
$\left[\mathcal{W}\Pi_{\Omega}\bm{s}_{\star}-\mathcal{W}\bm{s}\right]_{i}$ $=[\mathcal{W}\Pi_{\Omega}\bm{s}_{\star}]_{i}=\left[\mathcal{W}\Pi_{\Omega}(\bm{f}-\bm{z})-\mathcal{W}\Pi_{\Omega}(\bm{z}_{\star}-\bm{z})\right]_{i}$
$\leq|[\mathcal{W}\Pi_{\Omega}(\bm{f}-\bm{z})]_{i}|+|[\mathcal{W}\Pi_{\Omega}(\bm{z}_{\star}-\bm{z})]_{i}|$
$\leq 2\left\lVert[\mathcal{W}\Pi_{\Omega}(\bm{z}_{\star}-\bm{z})\right\lVert_{\infty}.$
Combing all the pieces we have $\big{\|}\mathcal{W}\Pi_{\Omega}\bm{s}_{\star}-\mathcal{W}\bm{s}\big{\|}_{\infty}\leq 2\left\lVert\mathcal{W}\Pi_{\Omega}(\bm{z}_{\star}-\bm{z})\right\lVert_{\infty}$.
Lemma 5.12.
(Non-expansiveness of $\mathcal{P}_{\mathcal{C}}$) For all $(\widetilde{{\bm{L}}},\widetilde{{\bm{R}}})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},C)$ with $C\geq(1+\varepsilon_{0})\sqrt{\mu c_{s}r/n}\sigma^{\star}_{1}$, $\varepsilon_{0}\in(0,1)$, and $\mathcal{P}_{\mathcal{C}}\begin{bmatrix}\widetilde{{\bm{L}}}\\ \widetilde{{\bm{R}}}\end{bmatrix}=\begin{bmatrix}{\bm{L}}\\ {\bm{R}}\end{bmatrix}$, we have
$\mathrm{dist}\left({\bm{L}},{\bm{R}};\bm{L}_{\star},\bm{R}_{\star}\right)\leq\mathrm{dist}\left(\widetilde{{\bm{L}}},\widetilde{{\bm{R}}};\bm{L}_{\star},\bm{R}_{\star}\right)$
and
$\max\{\|{\bm{L}}{\bm{R}}^{*}\|_{2,\infty},\|{\bm{R}}{\bm{L}}^{*}\|_{2,\infty}\}\leq C.$
Proof 5.13.
We first show the non-expansiveness of $\mathcal{P}_{\mathcal{C}}$. Let $\widetilde{{\bm{Q}}}\in\mathrm{GL}(r,\mathbb{C})$ be the best matrix that aligns $\left(\widetilde{{\bm{L}}},\widetilde{{\bm{R}}}\right)$ and $\left(\bm{L}_{\star},\bm{R}_{\star}\right)$. Notice that for $i\in[n_{1}]$ we have
$\big{\|}\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{\frac{1}{2}}\big{\|}_{2}$ $\leq\big{\|}\widetilde{{\bm{L}}}_{i,:}\widetilde{{\bm{Q}}}\bm{\Sigma}_{\star}^{1/2}\big{\|}_{2}\big{\|}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{\frac{1}{2}}\widetilde{{\bm{Q}}}^{-*}\bm{\Sigma}_{\star}^{-1/2}\big{\|}_{2}$
$\leq\big{\|}\widetilde{{\bm{L}}}_{i,:}\widetilde{{\bm{Q}}}\bm{\Sigma}_{\star}^{1/2}\big{\|}_{2}\big{\|}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{-1/2}\widetilde{{\bm{R}}}^{*}\big{\|}_{2}\big{\|}\widetilde{{\bm{R}}}\widetilde{{\bm{Q}}}^{-*}\bm{\Sigma}_{\star}^{-1/2}\big{\|}_{2}$
$\leq\big{\|}\widetilde{{\bm{L}}}_{i,:}\widetilde{{\bm{Q}}}\bm{\Sigma}_{\star}^{1/2}\big{\|}_{2}\left(\big{\|}(\widetilde{{\bm{R}}}\widetilde{{\bm{Q}}}^{-*}-\bm{R}_{\star})\bm{\Sigma}_{\star}^{-1/2}\big{\|}_{2}+\big{\|}\bm{R}_{\star}\bm{\Sigma}_{\star}^{-1/2}\big{\|}_{2}\right)$
$\leq(1+\varepsilon_{0})\big{\|}\widetilde{{\bm{L}}}_{i,:}\widetilde{{\bm{Q}}}\bm{\Sigma}_{\star}^{1/2}\big{\|}_{2},$
where the third inequality follows from $\big{\|}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{-1/2}\widetilde{{\bm{R}}}^{*}\big{\|}_{2}=1$, and the last inequality follows from
(15) $\left\lVert(\widetilde{{\bm{R}}}\widetilde{{\bm{Q}}}^{-*}-\bm{R}_{\star})\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}$ $\leq\left\lVert(\widetilde{{\bm{R}}}\widetilde{{\bm{Q}}}^{-*}-\bm{R}_{\star})\bm{\Sigma}_{\star}^{1/2}\bm{\Sigma}_{\star}^{-1}\right\lVert_{\mathrm{F}}$
(16) $\leq\left\lVert\bm{\Sigma}_{\star}^{-1}\right\lVert_{2}\mathrm{dist}\left(\widetilde{{\bm{L}}},\widetilde{{\bm{R}}};\bm{L}_{\star},\bm{R}_{\star}\right)\leq\varepsilon_{0}.$
Thus, as long as $C\geq(1+\varepsilon_{0})\sqrt{\mu c_{s}r/n}\sigma^{\star}_{1}$, $\forall i\in[n_{1}]$ we have
(17) $\frac{C}{\big{\|}\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{\frac{1}{2}}\big{\|}_{2}}\geq\frac{C}{(1+\varepsilon_{0})\big{\|}\widetilde{{\bm{L}}}_{i,:}\widetilde{{\bm{Q}}}\bm{\Sigma}_{\star}^{1/2}\big{\|}_{2}}\geq\frac{\big{\|}{\bm{L}}_{i,:}\bm{\Sigma}_{\star}^{1/2}\big{\|}_{2}}{\big{\|}\widetilde{{\bm{L}}}_{i,:}\widetilde{{\bm{Q}}}\bm{\Sigma}_{\star}^{1/2}\big{\|}_{2}},$
where the last inequality follows from $\bm{L}_{\star}\bm{\Sigma}_{\star}^{1/2}={\bm{U}}_{\star}\bm{\Sigma}_{\star}$ and
$\left\lVert[\bm{L}_{\star}]_{i,:}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}\leq\left\lVert{\bm{U}}_{\star}\right\lVert_{2,\infty}\left\lVert\bm{\Sigma}_{\star}\right\lVert_{2}\leq\sqrt{\mu c_{s}r/n}\sigma^{\star}_{1},\quad\forall i\in[n_{1}].$
For any vectors $\bm{u},\bm{v}\in\mathbb{R}^{n}$, it holds $\left\lVert\min(1,\lambda)\bm{u}-\bm{v}\right\lVert_{2}\leq\left\lVert\bm{u}-\bm{v}\right\lVert_{2}$ provided $\lambda\geq\frac{\left\lVert\bm{v}\right\lVert_{2}}{\left\lVert\bm{u}\right\lVert_{2}}$ (see [33, Claim 5]). Recalling that ${\bm{L}}_{i,:}=\min\left(1,\frac{C}{\big{\|}\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{\frac{1}{2}}\big{\|}_{2}}\right)\widetilde{{\bm{L}}}_{i,:}$, thus by Eq. 17 we have
$\left\lVert\left({\bm{L}}_{i,:}\widetilde{{\bm{Q}}}-[\bm{L}_{\star}]_{i,:}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2}\leq\left\lVert\left(\widetilde{{\bm{L}}}_{i,:}\widetilde{{\bm{Q}}}-[\bm{L}_{\star}]_{i,:}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2}$
and similarly $\left\lVert\left({\bm{R}}_{i,:}\widetilde{{\bm{Q}}}^{-*}-[\bm{R}_{\star}]_{i,:}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2}\leq\left\lVert\left(\widetilde{{\bm{R}}}_{i,:}\widetilde{{\bm{Q}}}^{-*}-[\bm{R}_{\star}]_{i,:}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2}$. Then, by definition of the metric we have the locally non-expensiveness of $\mathcal{P}_{\mathcal{C}}$, i.e.,
$\mathrm{dist}^{2}\left({\bm{L}},{\bm{R}};\bm{L}_{\star},\bm{R}_{\star}\right)$ $\leq\big{\|}\left({\bm{L}}\widetilde{{\bm{Q}}}-\bm{L}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\big{\|}_{\mathrm{F}}^{2}+\big{\|}\left({\bm{R}}\widetilde{{\bm{Q}}}^{-*}-\bm{R}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\big{\|}_{\mathrm{F}}^{2}$
$\leq\mathrm{dist}\left(\widetilde{{\bm{L}}},\widetilde{{\bm{R}}};\bm{L}_{\star},\bm{R}_{\star}\right).$
Noticing that $\|{\bm{L}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{1/2}\|_{2}=\|\min(1,\frac{C}{\|\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{\frac{1}{2}}\|_{2}})\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{1/2}\|_{2}\leq C$, $\forall i\in[n_{1}]$, thus $\|{\bm{L}}{\bm{R}}^{*}\|_{2,\infty}\leq\|{\bm{L}}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{1/2}\|_{2,\infty}\|{\bm{R}}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{-1/2}\|_{2}\leq C$. Similarly, we have $\|{\bm{R}}{\bm{L}}^{*}\|_{2,\infty}\leq C$.
5.2 Proof of Proposition 3.1
Proof 5.14.
For $d_{k}\leq\varepsilon_{0}\sigma^{\star}_{r}$, following Eq. 15, we have $\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}\leq\varepsilon_{0},$ thus
$\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}=\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-1/2}(\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2})^{*}\right\lVert_{\mathrm{F}}\leq\varepsilon_{0}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}.$
Similarly, we have $\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}\leq\varepsilon_{0}$ and $\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}\leq\varepsilon_{0}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}$. Therefore, we have
(18) $\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right\lVert_{\mathrm{F}}=$ $~{}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}+\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}+\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}$
(19) $\leq$ $~{}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}{\bm{V}}_{\star}^{*}\right\lVert_{\mathrm{F}}+\left\lVert{\bm{U}}_{\star}\bm{\Sigma}_{\star}^{1/2}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}$
(20) $\leq$ $~{}(1+\frac{\varepsilon_{0}}{2})\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right).$
Thus, by noticing $\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right)^{2}\leq 2d_{k}^{2}$, we complete the proof.
5.3 Proof of Theorem 3.3 (guaranteed initialization)
Proof 5.15.
This proof is under event (11) and we bound $d_{0}$ in the following. Following the notation in Algorithm 1, we denote
(21) $\bm{M}_{0}:=p^{-1}\mathcal{G}\left(\Pi_{\Omega}\bm{f}-\bm{s}_{0}\right),\quad\text{and}~{}\bm{M}_{0}^{(r)}~{}\text{be the top-$r$ SVD of}~{}\bm{M}_{0}.$
For some universal constant $\tilde{c}_{0}$, if provided $m\geq 64\tilde{c}_{0}^{2}\varepsilon_{0}^{-2}\kappa^{2}c_{s}\mu r^{2}\log n$, we have
(22) $\left\lVert\bm{M}_{0}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$ $\leq\left\lVert\bm{M}_{0}-p^{-1}\mathcal{G}\Pi_{\Omega}\bm{z}_{\star}\right\lVert_{2}+\left\lVert p^{-1}\mathcal{G}\Pi_{\Omega}\bm{z}_{\star}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$
(23) $=\left\lVert p^{-1}\mathcal{G}\left(\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\right)\right\lVert_{2}+\left\lVert\left(p^{-1}\mathcal{G}\Pi_{\Omega}-\mathcal{G}\right)\bm{z}_{\star}\right\lVert_{2}$
(24) $\leq 2\alpha n\big{\|}\mathcal{G}\big{(}\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\big{)}\big{\|}_{\infty}+\tilde{c}_{0}\sqrt{(\mu c_{s}r\log n)/m}\sigma^{\star}_{1}$
(25) $\leq 4\alpha n\big{\|}\mathcal{G}\bm{z}_{\star}\big{\|}_{\infty}+0.125\varepsilon_{0}\sigma^{\star}_{r}r^{-\frac{1}{2}}$
(26) $\leq 4\alpha c_{s}\mu r\sigma^{\star}_{1}+0.125\varepsilon_{0}\sigma^{\star}_{r}r^{-\frac{1}{2}},$
where the second step follows from $\|\mathcal{G}\left(\Pi_{\Omega}\bm{f}-\bm{s}_{0}\right)-\mathcal{G}\Pi_{\Omega}\bm{z}_{\star}\|_{2}=\|\mathcal{G}\left(\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\right)\|_{2}$. The third line follows from [2, Lemma 6] (as $\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}$ is at most $2\alpha m$-sparse) and Lemma 5.1. The fourth line follows from (14) with $\bm{z}=\bm{0}$, i.e., $\left\lVert\mathcal{G}\left(\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\right)\right\lVert_{\infty}=\left\lVert\mathcal{W}\left(\Pi_{\Omega}\bm{s}_{\star}-\bm{s}_{0}\right)\right\lVert_{\infty}\leq 2\left\lVert\mathcal{W}\Pi_{\Omega}\bm{z}_{\star}\right\lVert_{\infty}=2\left\lVert\mathcal{G}\Pi_{\Omega}\bm{z}_{\star}\right\lVert_{\infty}$. And the last inequality follows from
$\big{\|}\mathcal{G}\bm{z}_{\star}\big{\|}_{\infty}\leq\big{\|}{\bm{U}}_{\star}\big{\|}_{2,\infty}\big{\|}\mathcal{G}\bm{z}_{\star}\big{\|}_{2}\big{\|}{\bm{V}}_{\star}\big{\|}_{2,\infty}\leq\mu\frac{c_{s}r}{n}\left\lVert\mathcal{G}\bm{z}_{\star}\right\lVert_{2}.$
Therefore, we have
(27) $\left\lVert\bm{M}_{0}^{(r)}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$ $\leq\left\lVert\bm{M}_{0}^{(r)}-\bm{M}_{0}\right\lVert_{2}+\left\lVert\bm{M}_{0}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$
(28) $\leq 2\left\lVert\bm{M}_{0}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}~{}\leq 8\alpha c_{s}\mu r\kappa\sigma^{\star}_{r}+0.25\varepsilon_{0}\sigma^{\star}_{r}r^{-\frac{1}{2}},$
where the second inequality follows from the definition of $\bm{M}_{0}^{(r)}$ and Eckart-Young-Mirsky theorem. Then, by [33, Lemma 24] we have
$~{}\mathrm{dist}\left({\bm{U}}_{0}\bm{\Sigma}_{0}^{1/2},{\bm{V}}_{0}\bm{\Sigma}_{0}^{1/2};\bm{L}_{\star},\bm{R}_{\star}\right)\leq(\sqrt{2}+1)^{1/2}\left\lVert\bm{M}_{0}^{(r)}-\mathcal{G}\bm{z}_{\star}\right\lVert_{\mathrm{F}}$
$\leq$ $~{}\sqrt{2(\sqrt{2}+1)r}\left\lVert\bm{M}_{0}^{(r)}-\mathcal{G}\bm{z}_{\star}\right\lVert_{2}\leq 18\alpha c_{s}\kappa\mu r\sqrt{r}\sigma^{\star}_{r}+0.55\varepsilon_{0}\sigma^{\star}_{r}.$
Then, for any given $\varepsilon_{0}\in(0,1)$, we have $\mathrm{dist}\left({\bm{U}}_{0}\bm{\Sigma}_{0}^{1/2},{\bm{V}}_{0}\bm{\Sigma}_{0}^{1/2};\bm{L}_{\star},\bm{R}_{\star}\right)\leq\varepsilon_{0}\sigma^{\star}_{r}<\sigma^{\star}_{r}$ provided $0\leq\alpha\leq\frac{\varepsilon_{0}}{50c_{s}\kappa\mu r\sqrt{r}}$. Thus, for $c\geq 1+\varepsilon_{0}$, by Lemma 5.12 we have
$d_{0}\leq\mathrm{dist}\left({\bm{U}}_{0}\bm{\Sigma}_{0}^{1/2},{\bm{V}}_{0}\bm{\Sigma}_{0}^{1/2};\bm{L}_{\star},\bm{R}_{\star}\right)\leq\varepsilon_{0}\sigma^{\star}_{r},$
and $\max\{\left\lVert{\bm{L}}_{0}{\bm{R}}_{0}^{*}\right\lVert_{2,\infty},\left\lVert{\bm{R}}_{0}{\bm{L}}_{0}^{*}\right\lVert_{2,\infty}\}\leq c\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1}$. This completes the proof.
5.4 Proof of Theorems 3.5 and 3.8 (local convergence)
The proof follows a similar route established in [33], but the techniques and details are quite involved due to the extra Hankel structure, as well as the simultaneous presence of outliers and missing data.
In the following, we let $\varepsilon_{0}\in(0,1)$ be some fixed constant, and show that for any given $({\bm{L}}_{k},{\bm{R}}_{k})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},c\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1})$ we have the contraction $d_{k+1}\leq\tau d_{k}$ for some $\tau\in(0,1)$ provided $\varepsilon_{0}$ sufficiently small and $c\geq 1+\varepsilon_{0}$. We begin with some auxiliary lemmas.
Lemma 5.16.
For any $({\bm{L}}_{k},{\bm{R}}_{k})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},c\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1})$, we have
(29a) $\max\{\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2},\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}\}$ $\leq\varepsilon_{0},$
(29b) $\max\left\{\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty},\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}\right\}$ $\leq\frac{c}{1-\varepsilon_{0}}\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1},$
(29c) $\max\left\{\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty},\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\right\}$ $\leq\frac{c\kappa}{1-\varepsilon_{0}}\sqrt{\frac{c_{s}\mu r}{n}},$
(29d) $\left\lVert\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}$ $\leq\frac{1}{1-\varepsilon_{0}}.$
Proof 5.17.
The inequality (29a) follows from (15). To show (29b), noticing that
$\sigma_{r}(\bm{L}_{\natural}\bm{\Sigma}_{\star}^{-1/2})\geq\sigma_{r}(\bm{L}_{\star}\bm{\Sigma}_{\star}^{-1/2})-\sigma_{1}(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-1/2})\geq 1-\varepsilon_{0},$
then by $\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}\right\lVert_{2,\infty}\geq\sigma_{r}(\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2})\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}$ we have $\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}\leq\frac{\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}\right\lVert_{2,\infty}}{1-\varepsilon_{0}}$. Similarly, we have $\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}\leq\frac{\left\lVert{\bm{L}}_{k}{\bm{R}}_{k}^{*}\right\lVert_{2,\infty}}{1-\varepsilon_{0}}$ and conclude (29b). By using $\left\lVert{\bm{A}}{\bm{B}}\right\lVert_{2,\infty}\leq\left\lVert{\bm{A}}\right\lVert_{2,\infty}\left\lVert{\bm{B}}\right\lVert_{2}$, (29c) follows directly from (29b). (29d) follows from [33, Lemma 25] and (29a).
Lemma 5.18.
For all matrices $({\bm{L}}_{k},{\bm{R}}_{k})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},c\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1})$, it holds
$\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right),\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right)\right\rangle\right\rvert^{\frac{1}{2}}$
$\leq 2\sqrt{p}(1+\varepsilon_{0})\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right)$
provided $m\geq 6C_{\varepsilon_{0},\kappa}^{2}\varepsilon_{0}^{-2}c_{s}^{2}\mu^{2}r^{2}\log n$, where $C_{\varepsilon_{0},\kappa}=\left(1+\frac{c}{1-\varepsilon_{0}}\right)\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\kappa$.
Proof 5.19.
The decomposition ${\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}=\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}+\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}+\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}$ gives
$~{}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right),\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right)\right\rangle\right\rvert^{\frac{1}{2}}$
$\leq$ $~{}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}+\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right),\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}+\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\rangle\right\rvert^{\frac{1}{2}}$
$~{}+\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right),\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\rangle\right\rvert^{\frac{1}{2}}$
$\leq$ $~{}\sqrt{p\left(1+\varepsilon_{0}\right)}\left\lVert\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}{\bm{V}}_{\star}^{*}+{\bm{U}}_{\star}\bm{\Sigma}_{\star}^{1/2}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\lVert_{\mathrm{F}}+\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right),\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\rangle\right\rvert^{\frac{1}{2}}$
$\leq$ $~{}\sqrt{p\left(1+\varepsilon_{0}\right)}\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right)+\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right),\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\rangle\right\rvert^{\frac{1}{2}},$
where the first step follows from $\left\lvert\left\langle\Pi_{\Omega}(\bm{u}+\bm{v}),\bm{u}+\bm{v}\right\rangle\right\rvert^{\frac{1}{2}}\leq\left\lvert\left\langle\Pi_{\Omega}\bm{u},\bm{u}\right\rangle\right\rvert^{\frac{1}{2}}+\left\lvert\left\langle\Pi_{\Omega}\bm{v},\bm{v}\right\rangle\right\rvert^{\frac{1}{2}}$ for $\bm{u},\bm{v}\in\mathbb{C}^{m}$, the second inequality follows from Lemma 5.8, and in the last inequality we have used $\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}{\bm{V}}_{\star}^{*}\right\lVert_{\mathrm{F}}=\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}$. For the last term, provided $m\geq 6\varepsilon_{0}^{-2}C_{\varepsilon_{0},\kappa}^{2}c_{s}^{2}\mu^{2}r^{2}\log n$, we have
$~{}p^{-1}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right),\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\rangle\right\rvert$
$\leq$ $~{}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}^{2}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}$
$~{}+\sqrt{24p^{-1}n\log n}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{\mathrm{F}}$
$\leq$ $~{}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}^{2}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}+\varepsilon_{0}\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}\right)$
where the first inequality follows from Lemma 5.6, the second inequality follows from the inequalities $\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{\mathrm{F}}=\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\bm{\Sigma}_{\star}^{-1}\right\lVert_{\mathrm{F}}\leq\frac{1}{\sigma^{\star}_{r}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}$, and (by (29c))
(30) $\begin{split}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-\frac{1}{2}}\right\lVert_{2,\infty}&\leq\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{-\frac{1}{2}}\right\lVert_{2,\infty}+\left\lVert\bm{L}_{\star}\bm{\Sigma}_{\star}^{-\frac{1}{2}}\right\lVert_{2,\infty}\leq\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{\frac{c_{s}\mu r}{n}},\cr\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{\frac{1}{2}}\right\lVert_{2,\infty}&\leq\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{\frac{1}{2}}\right\lVert_{2,\infty}+\left\lVert\bm{L}_{\star}\bm{\Sigma}_{\star}^{\frac{1}{2}}\right\lVert_{2,\infty}\leq\left(1+\frac{c}{1-\varepsilon_{0}}\right)\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1}.\end{split}$
Therefore,
$~{}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right),\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\rangle\right\rvert^{\frac{1}{2}}$
$\leq$ $~{}\sqrt{p}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\sqrt{\varepsilon_{0}p}\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right).$
Similarly, we have
$~{}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right),\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\rangle\right\rvert^{\frac{1}{2}}$
$\leq$ $~{}\sqrt{p}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}+\sqrt{\varepsilon_{0}p}\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right).$
By (29a) and averaging the above two bounds, we conclude
$\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right),\mathcal{G}^{*}\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\rangle\right\rvert^{\frac{1}{2}}\leq 2\sqrt{p}\varepsilon_{0}\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right).$
Putting all the bound of terms together, we conclude the lemma.
Lemma 5.20.
For any indices set $\widehat{\Omega}\subseteq[n]$ we have
(31) $~{}\left\lvert\left\langle\Pi_{\widehat{\Omega}}\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right),\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right)\right\rangle\right\rvert^{\frac{1}{2}}$
(32) $\leq$ $~{}\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{c_{s}\mu r|\widehat{\Omega}|/(2n)}\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right),$
and for all ${\bm{L}}\in\mathbb{C}^{n_{1}\times r},~{}{\bm{R}}\in\mathbb{C}^{n_{2}\times r}$ it holds
$\left\lvert\left\langle\Pi_{\widehat{\Omega}}\mathcal{G}^{*}\left(\bm{L}\bm{R}^{*}\right),\mathcal{G}^{*}\left(\bm{L}\bm{R}^{*}\right)\right\rangle\right\rvert^{\frac{1}{2}}\leq\sqrt{|\widehat{\Omega}|}\min\{\left\lVert{\bm{L}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}\right\lVert_{\mathrm{F}},\left\lVert{\bm{R}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}\right\lVert_{\mathrm{F}}\}.$
Proof 5.21.
By using the decomposition ${\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}={\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}=(\bm{L}_{\natural}-\bm{L}_{\star})\bm{R}_{\natural}^{*}+\bm{L}_{\star}(\bm{R}_{\natural}-\bm{R}_{\star})=(\bm{L}_{\natural}-\bm{L}_{\star})\bm{R}_{\star}^{*}+\bm{L}_{\natural}(\bm{R}_{\natural}-\bm{R}_{\star})$, we have
$~{}2\left\lvert\left[\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\mathcal{G}\bm{z}_{\star}\right]_{i,j}\right\rvert$
$=$ $~{}\left\lvert\left[(\bm{L}_{\natural}-\bm{L}_{\star})\bm{R}_{\natural}^{*}+\bm{L}_{\star}(\bm{R}_{\natural}-\bm{R}_{\star})\right]_{i,j}+\left[(\bm{L}_{\natural}-\bm{L}_{\star})\bm{R}_{\star}^{*}+\bm{L}_{\natural}(\bm{R}_{\natural}-\bm{R}_{\star})\right]_{i,j}\right\rvert$
$\leq$ $~{}\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert[\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(i,:\right)\right\lVert_{2}+\left\lVert\bm{L}_{\star}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert[\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(j,:\right)\right\lVert_{2}$
$~{}+\left\lVert\bm{R}_{\star}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert[\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(i,:\right)\right\lVert_{2}+\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert[\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(j,:\right)\right\lVert_{2}$
$\leq$ $~{}\theta\left(\left\lVert[\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(i,:\right)\right\lVert_{2}+\left\lVert[\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(j,:\right)\right\lVert_{2}\right).$
where $\theta=\max\{\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}+\left\lVert\bm{R}_{\star}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty},\left\lVert\bm{L}_{\star}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}+\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\}\leq\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{c_{s}\mu r/n}$ due to Lemma 5.16. Let $\widehat{\Omega}=\{\hat{a}_{k}\}_{1\leq k\leq|\widehat{\Omega}|}$, denote $\widehat{\Phi}_{k}:=\{\left(i,j\right):\left(i+j-1\right)=\hat{a}_{k}\}$ and $\widehat{\Phi}=\cup_{1\leq k\leq|\widehat{\Omega}|}\widehat{\Phi}_{k}$. Then, we have
$~{}\left\lvert\left\langle\Pi_{\widehat{\Omega}}\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right),\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right)\right\rangle\right\rvert=\sum_{\hat{a}_{k}\in\widehat{\Omega}}\frac{1}{\varsigma_{\hat{a}_{k}}}\left\lvert\sum_{\left(i,j\right)\in\widehat{\Phi}_{k}}[\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\mathcal{G}\bm{z}_{\star}]_{i,j}\right\rvert^{2}$
$\leq$ $\sum_{\hat{a}_{k}\in\widehat{\Omega}}\sum_{\left(i,j\right)\in\widehat{\Phi}_{k}}\left\lvert[\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\mathcal{G}\bm{z}_{\star}]^{2}_{i,j}\right\rvert=\sum_{\left(i,j\right)\in\widehat{\Phi}}\left\lvert[\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\mathcal{G}\bm{z}_{\star}]^{2}_{i,j}\right\rvert$
$\leq$ $~{}\sum_{\left(i,j\right)\in\widehat{\Phi}}\frac{\theta^{2}}{4}\left(\left\lVert[\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(i,:\right)\right\lVert_{2}+\left\lVert[\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(j,:\right)\right\lVert_{2}\right)^{2}$
$\leq$ $~{}\frac{\theta^{2}}{2}\left(\sum_{j\in\widehat{\Phi}\left(:,i\right)}\sum_{i}\left\lVert[\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(i,:\right)\right\lVert_{2}^{2}+\sum_{i\in\widehat{\Phi}\left(:,j\right)}\sum_{j}\left\lVert[\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}]\left(j,:\right)\right\lVert_{2}^{2}\right)$
$\leq$ $~{}\frac{\theta^{2}}{2}\left(|\widehat{\Omega}|\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}+|\widehat{\Omega}|\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}\right),$
where the first inequality follows from $\left\lvert\widehat{\Phi}_{k}\right\rvert=\varsigma_{\hat{a}_{k}}$ and the Cauchy-Schwarz inequality, the third inequality follows from $(a+b)^{2}\leq 2a^{2}+2b^{2}$, and the fourth inequality follow from $\mathrm{card}\left(\widehat{\Phi}\left(i,:\right)\right)\leq|\widehat{\Omega}|$, $\mathrm{card}\left(\widehat{\Phi}\left(:,j\right)\right)\leq|\widehat{\Omega}|$ due to the Hankel structure. Therefore, by $\sqrt{a^{2}+b^{2}}\leq\left\lvert a\right\rvert+\left\lvert b\right\rvert$ we conclude the inequality (31). Similarly, by
$\left\lvert\left\langle\Pi_{\widehat{\Omega}}\mathcal{G}^{*}\left(\bm{L}\bm{R}^{*}\right),\mathcal{G}^{*}\left(\bm{L}\bm{R}^{*}\right)\right\rangle\right\rvert$ $\leq\sum_{\left(i,j\right)\in\widehat{\Phi}}\left\lvert[\bm{L}(i,:)\bm{R}^{*}(j,:)]^{2}\right\rvert\leq\sum_{\left(i,j\right)\in\widehat{\Phi}}\left\lVert\bm{L}(i,:)\right\lVert_{2}^{2}\left\lVert\bm{R}^{*}(j,:)\right\lVert_{2}^{2}$
$\leq|\widehat{\Omega}|\min\{\left\lVert{\bm{L}}\right\lVert_{2,\infty}^{2}\left\lVert{\bm{R}}\right\lVert_{\mathrm{F}}^{2},\left\lVert{\bm{R}}\right\lVert_{2,\infty}^{2}\left\lVert{\bm{L}}\right\lVert_{\mathrm{F}}^{2}\},$
we conclude the second statement.
Lemma 5.22.
Let $\bm{s}=\Gamma_{\gamma\alpha}\left(\Pi_{\Omega}\left(\bm{f}-\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}\right)\right)\right)$ with $\gamma>1$. Then, for all ${\bm{L}}_{{\bm{A}}}\in\mathbb{C}^{n_{1}\times r}$, ${\bm{R}}_{{\bm{A}}}\in\mathbb{C}^{n_{2}\times r}$, under event (12), (13) and provided $m\geq 6C_{\varepsilon_{0},\kappa}^{2}\varepsilon_{0}^{-2}c_{s}^{2}\mu^{2}r^{2}\log n$, we have
$~{}p^{-1}\left\lvert\left\langle\mathcal{G}\Pi_{\Omega}\left(\bm{s}-\bm{s}_{\star}\right),{\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right\rangle\right\rvert$
$\leq$ $~{}\sqrt{n}\xi\left(\left\lVert[\bm{\Delta}_{{\bm{L}}}\bm{\Sigma}_{\star}^{1/2}]\right\lVert_{\mathrm{F}}+\left\lVert[\bm{\Delta}_{{\bm{R}}}\bm{\Sigma}_{\star}^{1/2}]\right\lVert_{\mathrm{F}}\right)\min\{\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}},\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\},$
where $\xi=\left((\gamma+1)\alpha\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{\frac{c_{s}\mu r}{2}}+(2+2\varepsilon_{0})\frac{\sqrt{(\gamma+1)\alpha}}{\sqrt{(\gamma-1)}}\right).$
Proof 5.23.
Denote $\bm{\Delta}_{k}:={\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}$, and $\Omega^{s}:=\mathrm{supp}\left(\Pi_{\Omega}\bm{s}\right)$, $\Omega^{s_{\star}}:=\mathrm{supp}\left(\Pi_{\Omega}\bm{s}_{\star}\right)$. By the definition of $\bm{s}$, we have
(33) $\left\lvert\left[\Pi_{\Omega}\bm{s}_{\star}-\bm{s}\right]_{i}\right\rvert=\left\lvert\left[\Pi_{\Omega}\left(\bm{z}_{\star}-\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}\right)\right)\right]_{i}\right\rvert=\left\lvert\left[\Pi_{\Omega}(\mathcal{G}^{*}\bm{\Delta}_{k})\right]_{i}\right\rvert,~{}\text{for}~{}i\in\Omega^{s},$
where the last equation follows from $\mathcal{G}^{*}\mathcal{G}=\mathcal{I}$. For $i\in\Omega^{s_{\star}}\setminus\Omega^{s}$, by
$\bm{s}=\Gamma_{\gamma\alpha}\left(\Pi_{\Omega}\left(\bm{s}_{\star}-\mathcal{G}^{*}\left({\bm{L}}_{k}{\bm{R}}_{k}^{*}-\mathcal{G}\bm{z}_{\star}\right)\right)\right)=\Gamma_{\gamma\alpha}\left(\Pi_{\Omega}\left(\bm{s}_{\star}-\mathcal{G}^{*}\bm{\Delta}_{k}\right)\right),$
we know $\left\lvert\left[\Pi_{\Omega}\left(\bm{s}_{\star}-\mathcal{G}^{*}\bm{\Delta}_{k}\right)\right]_{i}\right\rvert$ is smaller than the $\gamma\alpha m$-th largest-in-magnitude entry of $\left\lvert\Pi_{\Omega}\left(\bm{s}_{\star}-\mathcal{G}^{*}\bm{\Delta}_{k}\right)\right\rvert$. Noticing $\Pi_{\Omega}\bm{s}_{\star}$ is at most $\alpha m$ sparse, thus $\left\lvert\left[\Pi_{\Omega}\left(\bm{s}_{\star}-\mathcal{G}^{*}\bm{\Delta}_{k}\right)\right]_{i}\right\rvert$ is smaller than the $(\gamma-1)\alpha m$-th largest entry of $\left\lvert\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right)\right\rvert$, and $\left\lvert\left[\Pi_{\Omega}\left(\bm{s}_{\star}-\mathcal{G}^{*}\bm{\Delta}_{k}\right)\right]_{i}\right\rvert^{2}\leq{\frac{\left\lvert\left\langle\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right),\mathcal{G}^{*}\bm{\Delta}_{k}\right\rangle\right\rvert}{(\gamma-1)\alpha m}}$. Then, we have
(34) $\left\lvert\left[\Pi_{\Omega}\bm{s}_{\star}\right]_{i}\right\rvert\leq{\frac{\left\lvert\left\langle\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right),\mathcal{G}^{*}\bm{\Delta}_{k}\right\rangle\right\rvert^{1/2}}{\sqrt{(\gamma-1)\alpha m}}}+\left\lvert\left[\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right)\right]_{i}\right\rvert,~{}\text{for}~{}i\in\Omega^{s_{\star}}\setminus\Omega^{s}.$
Hence,
$~{}\left\lvert\left\langle\mathcal{G}\Pi_{\Omega}\left(\bm{s}-\bm{s}_{\star}\right),{\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right\rangle\right\rvert=\left\lvert\left\langle\Pi_{\Omega}\left(\bm{s}-\bm{s}_{\star}\right),\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right\rangle\right\rvert$
$\leq$ $~{}\sum_{i\in\Omega^{s}}\left\lvert\left[\Pi_{\Omega}\left(\bm{s}-\bm{s}_{\star}\right)\right]_{i}\right\rvert\left\lvert\left[\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right]_{i}\right\rvert+\sum_{i\in\Omega^{s_{\star}}\setminus\Omega^{s}}\left\lvert\left[\Pi_{\Omega}\left(\bm{s}_{\star}\right)\right]_{i}\right\rvert\left\lvert\left[\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right]_{i}\right\rvert$
$\leq$ $~{}\sum_{i\in\Omega^{s}\cup\Omega^{s_{\star}}}\left\lvert\left[\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right)\right]_{i}\right\rvert\left\lvert\left[\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right]_{i}\right\rvert$
$~{}+\sum_{i\in\Omega^{s_{\star}}\setminus\Omega^{s}}{\frac{\left\lvert\left\langle\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right),\mathcal{G}^{*}\bm{\Delta}_{k}\right\rangle\right\rvert^{\frac{1}{2}}}{\sqrt{(\gamma-1)\alpha m}}\left\lvert\left[\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right]_{i}\right\rvert}$
$\leq$ $~{}\left\lvert\left\langle\Pi_{\Omega^{s}\cup\Omega^{s_{\star}}}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right),\mathcal{G}^{*}\bm{\Delta}_{k}\right\rangle\right\rvert^{\frac{1}{2}}\left\lvert\left\langle\Pi_{\Omega^{s}\cup\Omega^{s_{\star}}}\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}),\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right\rangle\right\rvert^{\frac{1}{2}}$
$~{}+\frac{\left\lvert\left\langle\Pi_{\Omega}\left(\mathcal{G}^{*}\bm{\Delta}_{k}\right),\mathcal{G}^{*}\bm{\Delta}_{k}\right\rangle\right\rvert}{2\beta(\gamma-1)}+\frac{\beta}{2}\left\lvert\left\langle\Pi_{\Omega^{s}\cup\Omega^{s_{\star}}}\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}),\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right\rangle\right\rvert$
$\leq$ $~{}(1+\frac{c\kappa}{1-\varepsilon_{0}})\sqrt{\frac{c_{s}\mu r}{2n}}(\gamma+1)\alpha m\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right)$
×$~{}\times\min\{\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}},\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\}+\frac{\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}(\bm{\Delta}_{k}),\mathcal{G}^{*}(\bm{\Delta}_{k})\right\rangle\right\rvert}{2\beta(\gamma-1)}$
$~{}+\frac{\beta(\gamma+1)\alpha m}{2}\min\{\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}^{2}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2},\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}^{2}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}\},$
where the third inequality follows form $ab\leq\frac{\beta}{2}a^{2}+\frac{b^{2}}{2\beta}$ for any $\beta>0$, and the Cauchy-Schwarz inequality, the fourth inequality follows from Lemma 5.20. Letting
$\beta=\frac{\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}(\bm{\Delta}_{k}),\mathcal{G}^{*}(\bm{\Delta}_{k})\right\rangle\right\rvert^{\frac{1}{2}}}{(\sqrt{(\gamma-1)(\gamma+1)\alpha m}\min\{\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}},\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\})},$
we then have
$~{}\left\lvert\left\langle\mathcal{G}\Pi_{\Omega}\left(\bm{s}-\bm{s}_{\star}\right),{\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right\rangle\right\rvert$
$\leq$ $~{}\Big{[}\Big{(}1+\frac{c\kappa}{1-\varepsilon_{0}}\Big{)}\sqrt{\frac{c_{s}\mu r}{2n}}(\gamma+1)\alpha m\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right)+$
$~{}\frac{\sqrt{(\gamma+1)\alpha m}}{\sqrt{(\gamma-1)}}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}(\bm{\Delta}_{k}),\mathcal{G}^{*}(\bm{\Delta}_{k})\right\rangle\right\rvert^{\frac{1}{2}}\Big{]}\min\{\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}},\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\}$
$\leq$ $~{}p\sqrt{n}\xi\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right)\min\{\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}},\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\},$
where $\xi=\left((\gamma+1)\alpha\left(1+\frac{c\kappa}{1-\varepsilon_{0}}\right)\sqrt{\frac{c_{s}\mu r}{2}}+(2+2\varepsilon_{0})\frac{\sqrt{(\gamma+1)\alpha}}{\sqrt{(\gamma-1)}}\right)$ and the last inequality follows from Lemma 5.18.
5.4.1 Contraction Property in Full Observation
Proof 5.24 (Proof of Theorem 3.5).
For the case of full observation $m=n$ ($p=1$), we show the contraction property for one step update of ${{\bm{L}}}_{k}$, ${{\bm{R}}}_{k}$ in terms of the metric $d_{k}$. Denote
$\widetilde{{\bm{L}}}_{k+1}$ $:={\bm{L}}_{k}-\eta\nabla_{{\bm{L}}}\ell({\bm{L}}_{k},{\bm{R}}_{k};\bm{s}_{k+1})\left({\bm{R}}^{*}_{k}{\bm{R}}_{k}\right)^{-1}$
$={\bm{L}}_{k}-\eta\mathcal{G}\left(\bm{s}_{k+1}-\bm{f}\right){\bm{R}}_{k}\left({\bm{R}}^{*}_{k}{\bm{R}}_{k}\right)^{-1}-\eta{\bm{L}}_{k}.$
Then, we have
$~{}\left(\widetilde{{\bm{L}}}_{k+1}{\bm{Q}}_{k}-\bm{L}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}=\left[\bm{\Delta}_{{\bm{L}}_{k}}-\eta\mathcal{G}\big{[}\bm{s}_{k+1}-\bm{z}_{\star}-\bm{s}_{\star}\big{]}\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}-\eta\bm{L}_{\natural}\right]\bm{\Sigma}_{\star}^{1/2}$
$=$ $~{}\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}-\eta\big{[}\mathcal{G}\left(\bm{s}_{k+1}-\bm{s}_{\star}\right)+\left(\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\bm{L}_{\star}\bm{R}_{\star}^{*}\right)\big{]}\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}$
$=$ $~{}\underbrace{(1-\eta)\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}-\eta\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}}_{\bm{T}_{1}}$
$~{}-\underbrace{\eta[\mathcal{G}\left(\bm{s}_{k+1}-\bm{s}_{\star}\right)]\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}}_{\bm{T}_{2}},$
where in the third equation we have used the decomposition $\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\bm{L}_{\star}\bm{R}_{\star}^{*}=\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\natural}^{*}+\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}=\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}+\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}+\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}$. Noticing that
$\left\lVert\left(\widetilde{{\bm{L}}}_{k+1}{\bm{Q}}_{k}-\bm{L}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}=\left\lVert\bm{T}_{1}\right\lVert_{\mathrm{F}}^{2}+\left\lVert\bm{T}_{2}\right\lVert_{\mathrm{F}}^{2}-2\mathrm{Re}\left\langle\bm{T}_{1},\bm{T}_{2}\right\rangle,$
and we have the following claims, whose proof is deferred later in this subsection:
Claim 1.
For $({\bm{L}}_{k},{\bm{R}}_{k})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},c\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1})$, we have
$\left\lVert\bm{T}_{1}\right\lVert_{\mathrm{F}}^{2}\leq$ $~{}\left((1-\eta)^{2}+\frac{2\varepsilon_{0}}{1-\varepsilon_{0}}\eta(1-\eta)+\frac{2\varepsilon_{0}+\varepsilon_{0}^{2}}{(1-\varepsilon_{0})^{2}}\eta^{2}\right)d_{k}^{2},$
$\left\lvert\left\langle\bm{T}_{1},\bm{T}_{2}\right\rangle\right\rvert\leq$ $~{}\frac{(\eta-\eta^{2})c\kappa\sqrt{c_{s}\mu r}\xi}{2(1-\varepsilon_{0})^{3}}\left(3\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}\right)$
$~{}+\frac{\eta^{2}\sqrt{c_{s}\mu r}\xi}{2(1-\varepsilon_{0})^{2}}\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}+3\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}\right),$
$\left\lVert\bm{T}_{2}\right\lVert_{\mathrm{F}}\leq$ $~{}\frac{2\eta c\kappa\sqrt{c_{s}\mu r}\xi}{(1-\varepsilon_{0})^{3}}\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right).$
Therefore, if provided $\alpha\leq\frac{\varepsilon^{2}_{0}(\gamma-1)}{64(\gamma+1)(1+c\kappa)^{2}c_{s}\mu r}$, combing all the pieces above we have
$\quad\left\lVert\left(\widetilde{{\bm{L}}}_{k+1}{\bm{Q}}_{k}-\bm{L}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}$
$\leq C_{1}(\varepsilon_{0},\eta)d_{k}^{2}+C_{2}(\varepsilon_{0},\eta)\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}+C_{3}(\varepsilon_{0},\eta)\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}$
$\leq\left(1-0.6\eta\right)^{2}d_{k}^{2},\quad\text{for}~{}\varepsilon_{0}=0.01,~{}0<\eta\leq 1,$
where we have used the Cauchy-Schwarz inequality and and the constants satisfy $C_{1}(\varepsilon_{0},\eta)\leq(1-\eta)^{2}+\frac{2\varepsilon_{0}}{1-\varepsilon_{0}}\eta(1-\eta)+\frac{8\eta^{2}\varepsilon_{0}^{2}}{(1-\varepsilon_{0})^{6}}$, $C_{2}(\varepsilon_{0},\eta)\leq\frac{3\eta\varepsilon_{0}-2\eta^{2}\varepsilon_{0}-\eta^{2}\varepsilon_{0}^{2}}{2(1-\varepsilon_{0})^{3}}$, $C_{3}(\varepsilon_{0},\eta)\leq\frac{\eta\varepsilon_{0}+2\eta^{2}\varepsilon_{0}-3\eta^{2}\varepsilon_{0}^{2}}{2(1-\varepsilon_{0})^{3}}$. Similarly, let $\widetilde{{\bm{R}}}_{k+1}:={\bm{R}}_{k}-\eta\nabla_{{\bm{R}}}\ell({\bm{L}}_{k},{\bm{R}}_{k};\bm{s}_{k+1})\left({\bm{R}}^{*}_{k}{\bm{R}}_{k}\right)^{-1},$ we can show $\left\lVert\left(\widetilde{{\bm{R}}}_{k+1}{\bm{Q}}_{k}^{-*}-\bm{R}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}\leq(1-0.6\eta)^{2}d_{k}^{2}$. Thus we have
$\mathrm{dist}\left(\widetilde{{\bm{L}}}_{k+1},\widetilde{{\bm{R}}}_{k+1};\bm{L}_{\star},\bm{R}_{\star}\right)\leq\left(1-0.6\eta\right)d_{k}.$
Then, Lemma 5.12 gives $d_{k+1}\leq(1-0.6\eta)d_{k}$ and $({\bm{L}}_{k+1},{\bm{R}}_{k+1})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},c\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1})$. Finally, we complete the proof by proving Claim 1.
Proof 5.25 (Proof of Claim 1).
The upper bound of $\left\lVert\bm{T}_{1}\right\lVert_{\mathrm{F}}$ following the argument as in [33, Eq. (46)]. For $\left\lvert\left\langle\bm{T}_{1},\bm{T}_{2}\right\rangle\right\rvert$, we denote
(35) $\begin{split}{\bm{L}}_{{\bm{A}}_{1}}&:=\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2},~{}{\bm{R}}_{{\bm{A}}_{1}}:=\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2},\cr{\bm{L}}_{{\bm{A}}_{2}}&:=\bm{L}_{\star}\bm{\Sigma}_{\star}^{-1/2}={\bm{U}}_{\star},~{}{\bm{R}}_{{\bm{A}}_{2}}:=\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{R}_{\natural}^{*}\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}.\end{split}$
Then, by applying Lemma 5.22 with $m=n$ we have
$~{}\left\lvert\left\langle\bm{T}_{1},\bm{T}_{2}\right\rangle\right\rvert=\Big{\lvert}\eta(1-\eta)\mathrm{tr}\left([\mathcal{G}\left(\bm{s}_{k+1}-\bm{s}_{\star}\right)]\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}\bm{\Delta}_{{\bm{L}}_{k}}^{*}\right)$
$~{}\qquad\quad-\eta^{2}\mathrm{tr}\left([\mathcal{G}\left(\bm{s}_{k+1}-\bm{s}_{\star}\right)]\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{R}_{\natural}^{*}\bm{\Delta}_{{\bm{R}}_{k}}\bm{L}_{\star}^{*}\right)\Big{\rvert}$
$\leq$ $~{}\eta(1-\eta)\left\lvert\left\langle\mathcal{G}\left(\bm{s}_{k+1}-\bm{s}_{\star}\right),{\bm{L}}_{{\bm{A}}_{1}}{\bm{R}}_{{\bm{A}}_{1}}^{*}\right\rangle\right\rvert+\eta^{2}\left\lvert\left\langle\mathcal{G}\left(\bm{s}_{k+1}-\bm{s}_{\star}\right),{\bm{L}}_{{\bm{A}}_{2}}{\bm{R}}_{{\bm{A}}_{2}}^{*}\right\rangle\right\rvert$
$\leq$ $~{}\sqrt{n}\xi\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right)\times$
$~{}\left((\eta-\eta^{2})\left\lVert{\bm{R}}_{{\bm{A}}_{1}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}_{1}}\right\lVert_{\mathrm{F}}+\eta^{2}\left\lVert{\bm{L}}_{{\bm{A}}_{2}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}_{2}}\right\lVert_{\mathrm{F}}\right),$
together with $ab\leq\frac{a^{2}+b^{2}}{2}$, (29c), (29d) and the fact that
(36) $\left\lVert\bm{\Sigma}_{\star}^{1/2}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}$ $=\left\lVert\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2},$
(37) $\left\lVert\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{R}_{\natural}^{*}\right\lVert_{2}$ $=\left\lVert\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2},$
(38) (29d)$\left\lVert{\bm{L}}_{{\bm{A}}_{1}}\right\lVert_{\mathrm{F}}\leq\frac{1}{(1-\varepsilon_{0})^{2}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}},$ $~{}\left\lVert{\bm{R}}_{{\bm{A}}_{2}}\right\lVert_{\mathrm{F}}\leq\frac{1}{(1-\varepsilon_{0})^{2}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}},$
we conclude the second inequality in Claim 1. For $\left\lVert\bm{T}_{2}\right\lVert_{\mathrm{F}}$, using the variational representation of the Frobenius norm, for some ${\bm{L}}_{{\bm{A}}}\in\mathbb{C}^{n_{1}\times r}$, $\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}=1$ we have
$\quad\left\lVert\bm{T}_{2}\right\lVert_{\mathrm{F}}=\eta\left\lvert\left\langle[\mathcal{G}\left(\bm{s}_{k+1}-\bm{s}_{\star}\right)]\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2},{\bm{L}}_{{\bm{A}}}\right\rangle\right\rvert$
$=\eta\left\lvert\left\langle[\mathcal{G}\left(\bm{s}_{k+1}-\bm{s}_{\star}\right)],{\bm{L}}_{{\bm{A}}}\bm{\Sigma}_{\star}^{1/2}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}(\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2})^{*}\right\rangle\right\rvert$
$\leq\eta\sqrt{n}\xi\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right)\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-\frac{1}{2}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\bm{\Sigma}_{\star}^{1/2}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}$
$\leq\frac{2\eta c\kappa\sqrt{c_{s}\mu r}\xi}{(1-\varepsilon_{0})^{3}}\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right),$
where the first inequality follows from Lemma 5.22, the last inequality follows from (29c) and similar argument to (36).
5.4.2 Contraction Property in Partial Observation
The following lemma is crucial for the proof of the theorem.
Lemma 5.26.
For any ${\bm{L}}_{{\bm{A}}},{\bm{L}}_{{\bm{B}}}\in\mathbb{C}^{n_{1}\times r}$, ${\bm{R}}_{{\bm{A}}},{\bm{R}}_{{\bm{B}}}\in\mathbb{C}^{n_{2}\times r}$ independent of $\Omega$, with probability at least $1-2n^{-2}$ it holds
$\left\lvert\left\langle\mathcal{G}\left(p^{-1}\Pi_{\Omega}-\mathcal{I}\right)\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right),{\bm{L}}_{{\bm{B}}}{\bm{R}}_{{\bm{B}}}^{*}\right\rangle\right\rvert$ $\leq\sqrt{24p^{-1}n\log n}\min\{\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{B}}}\right\lVert_{\mathrm{F}},$
$\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{L}}_{{\bm{B}}}\right\lVert_{2,\infty}\}$ $\min\{\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{B}}}\right\lVert_{\mathrm{F}},\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{B}}}\right\lVert_{2,\infty}\}$
provided $m\geq 24\log n$.
Proof 5.27.
The proof is deferred to Appendix A.
Proof 5.28 (Proof of Theorem 3.8).
Similar to the proof of Theorem 3.5, let $\widetilde{{\bm{L}}}_{k+1}:={\bm{L}}_{k}-\eta\nabla_{{\bm{L}}}\ell({\bm{L}}_{k},{\bm{R}}_{k};\bm{s}_{k+1})\left({\bm{R}}^{*}_{k}{\bm{R}}_{k}\right)^{-1}$ we have
$\left(\widetilde{{\bm{L}}}_{k+1}{\bm{Q}}_{k}-\bm{L}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}$
$=$ $\left[\bm{\Delta}_{{\bm{L}}_{k}}-\eta\mathcal{G}\big{[}p^{-1}\Pi_{\Omega}\left(\mathcal{G}^{*}\left(\bm{L}_{\natural}\bm{R}_{\natural}^{*}\right)+\bm{s}_{k+1}-\bm{f}\right)-\mathcal{G}^{*}(\bm{L}_{\natural}\bm{R}_{\natural}^{*})\big{]}\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}-\eta\bm{L}_{\natural}\right]\bm{\Sigma}_{\star}^{1/2}$
$=$ $\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}-\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\bm{L}_{\star}\bm{R}_{\star}^{*}\right)\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}$
$~{}-\eta p^{-1}[\mathcal{G}\Pi_{\Omega}\left(\bm{s}_{k+1}-\bm{s}_{\star}\right)]\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}-\eta\left(\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\bm{L}_{\star}\bm{R}_{\star}^{*}\right)\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}$
$=$ $(1-\eta)\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}-\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}+\bm{L}_{\natural}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}$
$~{}-\eta p^{-1}[\mathcal{G}\Pi_{\Omega}\left(\bm{s}_{k+1}-\bm{s}_{\star}\right)]\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}-\eta\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}$
$=$ $\underbrace{(1-\eta)\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}-\eta\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}}_{\bm{T}_{1}}$
$~{}-\underbrace{\eta p^{-1}[\mathcal{G}\Pi_{\Omega}\left(\bm{s}_{k+1}-\bm{s}_{\star}\right)]\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}}_{\bm{T}_{3}}$
$~{}-\underbrace{\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}+\bm{L}_{\natural}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}}_{\bm{T}_{4}},$
where the second equation follows from $\bm{f}=\bm{z}_{\star}+\bm{s}_{\star}$, $\mathcal{G}^{*}(\bm{L}_{\star}\bm{R}_{\star}^{*})=\bm{z}_{\star}$ and $\mathcal{G}\mathcal{G}^{*}(\bm{L}_{\star}\bm{R}_{\star}^{*})=\mathcal{G}\bm{z}_{\star}=\bm{L}_{\star}\bm{R}_{\star}^{*}$, the third equation follows from $\bm{L}_{\natural}\bm{R}_{\natural}^{*}-\bm{L}_{\star}\bm{R}_{\star}^{*}=\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\natural}^{*}+\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}=\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}+\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}+\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}$. Noticing that
(39) $~{}\left\lVert\left(\widetilde{{\bm{L}}}_{k+1}{\bm{Q}}_{k}-\bm{L}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}$
(40) $=$ $~{}\left\lVert\bm{T}_{1}\right\lVert_{\mathrm{F}}^{2}+\left\lVert\bm{T}_{3}\right\lVert_{\mathrm{F}}^{2}+\left\lVert\bm{T}_{4}\right\lVert_{\mathrm{F}}^{2}-2\mathrm{Re}\left\langle\bm{T}_{1},\bm{T}_{3}\right\rangle-2\mathrm{Re}\left\langle\bm{T}_{1},\bm{T}_{4}\right\rangle+2\mathrm{Re}\left\langle\bm{T}_{3},\bm{T}_{4}\right\rangle,$
thus we bound $\left\lVert\left(\widetilde{{\bm{L}}}_{k+1}{\bm{Q}}_{k}-\bm{L}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}$ term by term by the following claims:
Claim 2.
For $({\bm{L}}_{k},{\bm{R}}_{k})\in\mathcal{E}(\varepsilon_{0}\sigma^{\star}_{r},c\sqrt{\frac{c_{s}\mu r}{n}}\sigma^{\star}_{1})$, $c\geq 1+\varepsilon_{0}$, $\varepsilon_{0},\eta\in(0,1)$, if provided $m\geq 24c^{2}(1+c\kappa)^{2}\varepsilon_{0}^{-2}c_{s}^{2}\mu^{2}r^{2}\log n$, we have
$\left\lvert\left\langle\bm{T}_{1},\bm{T}_{3}\right\rangle\right\rvert$ $\leq\frac{(\eta-\eta^{2})c\kappa\sqrt{c_{s}\mu r}\xi}{2(1-\varepsilon_{0})^{3}}\left(3\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}\right)$
$\quad~{}+\frac{\eta^{2}\sqrt{c_{s}\mu r}\xi}{2(1-\varepsilon_{0})^{2}}\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}+3\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}\right),$
$\left\lVert\bm{T}_{3}\right\lVert_{\mathrm{F}}$ $\leq\frac{2\eta c\kappa\sqrt{c_{s}\mu r}\xi}{(1-\varepsilon_{0})^{3}}\left(\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\right),$
$\left\lvert\left\langle\bm{T}_{1},\bm{T}_{4}\right\rangle\right\rvert$ $\leq\left(\frac{(1-\eta)\eta\varepsilon_{0}}{1-\varepsilon_{0}}+\frac{\varepsilon_{0}\eta(1+\varepsilon_{0})}{(1-\varepsilon_{0})^{4}}\right)\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}$
$\quad~{}+\left(\frac{\varepsilon_{0}\eta^{2}}{(1-\varepsilon_{0})^{2}}+\frac{\varepsilon_{0}\eta(1+\varepsilon_{0})}{(1-\varepsilon_{0})^{4}}\right)\left\lVert\bm{\Sigma}_{\star}^{1/2}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}^{2},$
$\left\lVert\bm{T}_{4}\right\lVert_{\mathrm{F}}$ $\leq\frac{\varepsilon_{0}\eta}{(1-\varepsilon_{0})^{2}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\frac{2\eta(1+c\kappa)}{(1-\varepsilon_{0})^{3}}\sqrt{\frac{24c_{s}^{2}\mu^{2}r^{2}\log n}{m}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}.$
Therefore, for $c\geq 1+\varepsilon_{0}$, $\varepsilon_{0},\eta\in(0,1)$, if provided $m\geq 24c^{2}(1+c\kappa)^{2}\varepsilon_{0}^{-2}c_{s}^{2}\mu^{2}r^{2}\log n$ and $\alpha\leq\frac{\varepsilon^{2}_{0}(\gamma-1)}{64(\gamma+1)(1+c\kappa)^{2}c_{s}\mu r}$, by combing all the pieces above together with $\left\lvert\left\langle\bm{T}_{3},\bm{T}_{4}\right\rangle\right\rvert\leq\left\lVert\bm{T}_{3}\right\lVert_{\mathrm{F}}\left\lVert\bm{T}_{4}\right\lVert_{\mathrm{F}}$ and (39) we have
$\left\lVert\left(\widetilde{{\bm{L}}}_{k+1}{\bm{Q}}_{k}-\bm{L}_{\star}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}$ $\leq C_{1}(\varepsilon_{0},\eta)d_{k}^{2}+C_{2}(\varepsilon_{0},\eta)\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}+C_{3}(\varepsilon_{0},\eta)\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}$
$\leq\left(1-0.6\eta\right)^{2}d_{k}^{2},\quad\text{for}~{}\varepsilon_{0}=0.01,~{}0<\eta\leq 0.6,$
where in the first inequality we have used the Cauchy-Schwarz inequality, and the constants satisfy $C_{1}(\varepsilon_{0},\eta)\leq(1-\eta)^{2}+\frac{2\varepsilon_{0}}{1-\varepsilon_{0}}\eta(1-\eta)+\frac{2\varepsilon_{0}+\varepsilon_{0}^{2}}{(1-\varepsilon_{0})^{2}}\eta^{2}~{}+\frac{8\eta^{2}\varepsilon_{0}^{2}}{(1-\varepsilon_{0})^{8}}$, $C_{2}(\varepsilon_{0},\eta)\leq\frac{\varepsilon_{0}^{2}\eta^{2}+8\varepsilon_{0}\eta+2\varepsilon_{0}^{2}\eta-5\varepsilon_{0}\eta^{2}}{(1-\varepsilon_{0})^{4}}+\frac{2(1-\eta)\eta\varepsilon_{0}}{1-\varepsilon_{0}}+\frac{8\eta^{2}\varepsilon_{0}^{2}}{(1-\varepsilon_{0})^{7}}$, $C_{3}(\varepsilon_{0},\eta)\leq\frac{4\eta\varepsilon_{0}+\eta^{2}\varepsilon_{0}+2\eta\varepsilon_{0}^{2}-3\eta^{2}\varepsilon_{0}^{2}}{(1-\varepsilon_{0})^{4}}+\frac{2\varepsilon_{0}\eta^{2}}{(1-\varepsilon_{0})^{2}}+\frac{12\eta^{2}\varepsilon_{0}^{2}}{(1-\varepsilon_{0})^{7}}$. Similarly, let $\widetilde{{\bm{R}}}_{k+1}:={\bm{R}}_{k}-\eta\nabla_{{\bm{R}}}\ell({\bm{L}}_{k},{\bm{R}}_{k};\bm{s}_{k+1})\left({\bm{R}}^{*}_{k}{\bm{R}}_{k}\right)^{-1},$ we can use the same strategy to show $\|(\widetilde{{\bm{R}}}_{k+1}{\bm{Q}}_{k}^{-*}-\bm{R}_{\star})\bm{\Sigma}_{\star}^{1/2}\|_{\mathrm{F}}^{2}\leq(1-0.6\eta)^{2}d_{k}^{2}$. Thus we have
$\mathrm{dist}\left(\widetilde{{\bm{L}}}_{k+1},\widetilde{{\bm{R}}}_{k+1};\bm{L}_{\star},\bm{R}_{\star}\right)\leq\left(1-0.6\eta\right)d_{k}.$
By Lemma 5.12, we conclude the statement in Theorem 3.8. Finally, we finish the proof by proving Claim 2.
Proof 5.29 (Proof of Claim 2).
The proof of the upper bound for $\left\lvert\left\langle\bm{T}_{1},\bm{T}_{3}\right\rangle\right\rvert$ and $\left\lVert\bm{T}_{3}\right\lVert_{\mathrm{F}}$ are similar to $\left\lvert\left\langle\bm{T}_{1},\bm{T}_{2}\right\rangle\right\rvert$ and $\left\lVert\bm{T}_{2}\right\lVert_{\mathrm{F}}$ as in Claim 1, respectively. For $\left\lvert\left\langle\bm{T}_{1},\bm{T}_{4}\right\rangle\right\rvert$, we have:
$~{}\left\langle\bm{T}_{1},\bm{T}_{4}\right\rangle=\left\langle(1-\eta)\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2},\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}\right)\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\rangle$
$~{}+\left\langle(1-\eta)\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2},\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{L}_{\natural}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\rangle$
$~{}-\left\langle\eta\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2},\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}\right)\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\rangle$
$~{}-\left\langle\eta\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2},\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{L}_{\natural}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\rangle$
Noticing $\bm{R}_{\natural}=\bm{R}_{\star}+\bm{\Delta}_{{\bm{R}}_{k}}$, let
(41) $\begin{split}{\bm{L}}_{{\bm{A}}_{3}}&:=\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2},~{}{\bm{R}}_{{\bm{A}}_{3}}:=\bm{R}_{\star}\bm{\Sigma}_{\star}^{-1/2},\cr{\bm{L}}_{{\bm{A}}_{4}}&:=\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2},~{}{\bm{R}}_{{\bm{A}}_{4}}:=\bm{\Delta}_{{\bm{R}}_{k}}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2},\end{split}$
by Lemmas 5.6 and 5.26 we then have
$~{}\left\lvert\left\langle(1-\eta)\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2},\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}\right)\left(\bm{R}_{\star}+\bm{\Delta}_{{\bm{R}}_{k}}\right)\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\rangle\right\rvert$
$\leq$ $~{}(1-\eta)\eta\varepsilon_{0}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{R}_{\star}^{*}\right\lVert_{\mathrm{F}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}\right\lVert_{\mathrm{F}}$
$~{}+(1-\eta)\eta\left\lvert\left\langle\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}(\bm{R}_{\star}\bm{\Sigma}_{\star}^{-1/2})^{*}\right),\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\rangle\right\rvert$
$\leq$ $~{}\frac{(1-\eta)\eta\varepsilon_{0}}{1-\varepsilon_{0}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}+(1-\eta)\eta\sqrt{24p^{-1}n\log n}\left\lVert{\bm{L}}_{{\bm{A}}_{3}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{A}}_{3}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}_{4}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}_{4}}\right\lVert_{\mathrm{F}}$
$\leq$ $~{}\frac{(1-\eta)\eta\varepsilon_{0}}{1-\varepsilon_{0}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{\frac{1}{2}}\right\lVert_{\mathrm{F}}^{2}+\frac{(1+c)(\eta-\eta^{2})}{(1-\varepsilon_{0})^{3}}\sqrt{\frac{24c_{s}^{2}\mu^{2}\kappa^{2}r^{2}\log n}{m}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{\frac{1}{2}}\right\lVert_{\mathrm{F}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{\frac{1}{2}}\right\lVert_{\mathrm{F}},$
where the last step follows from (29d), (30) and $\left\lVert{\bm{R}}_{{\bm{A}}_{3}}\right\lVert_{2,\infty}=\left\lVert{\bm{U}}_{\star}\right\lVert_{2,\infty}$, $\left\lVert{\bm{R}}_{{\bm{A}}_{4}}\right\lVert_{\mathrm{F}}\leq\|\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\|_{\mathrm{F}}\|\bm{\Sigma}_{\star}^{-1}\|_{2}\|\bm{\Sigma}_{\star}^{1/2}(\bm{R}_{\natural}^{*}\bm{R}_{\natural})^{-1}\bm{\Sigma}_{\star}^{1/2}\|_{2}\leq\frac{1}{\sigma^{\star}_{r}(1-\varepsilon_{0})^{2}}\|\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\|_{\mathrm{F}}$. Similarly, by Lemma 5.26 and the definition of ${\bm{L}}_{{\bm{A}}_{2}}$, ${\bm{L}}_{{\bm{A}}_{3}}$, ${\bm{R}}_{{\bm{A}}_{2}}$, ${\bm{R}}_{{\bm{A}}_{5}}$ in (35) and (41), we have
$~{}\left\lvert\left\langle\eta\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2},\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}\right)\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\rangle\right\rvert$
$=$ $~{}\left\lvert\left\langle\eta(\bm{L}_{\star}\bm{\Sigma}_{\star}^{-\frac{1}{2}})\bm{\Sigma}_{\star}^{\frac{1}{2}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{R}_{\natural}^{*},\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}\right)\right\rangle\right\rvert$
$\leq$ $~{}\eta^{2}\sqrt{24p^{-1}n\log n}\left\lVert{\bm{L}}_{{\bm{A}}_{2}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}_{3}}\right\lVert_{\mathrm{F}}\left\lVert\bm{R}_{\star}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}_{2}}\right\lVert_{\mathrm{F}}$
$\leq$ $~{}\frac{\eta^{2}}{(1-\varepsilon_{0})^{2}}\sqrt{\frac{24c_{s}^{2}\mu^{2}r^{2}\log n}{m}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}},$
and
$~{}\left\lvert\left\langle(1-\eta)\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2},\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{L}_{\natural}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\rangle\right\rvert$
$=$ $~{}\left\lvert\left\langle(1-\eta)\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}(\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2})^{*},\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{L}_{\natural}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\rangle\right\rvert$
$\leq$ $~{}(1-\eta)\eta\sqrt{24p^{-1}n\log n}\left\lVert{\bm{L}}_{{\bm{A}}_{1}}\right\lVert_{\mathrm{F}}\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert\bm{R}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}$
$\leq$ $~{}\frac{(1-\eta)\eta}{(1-\varepsilon_{0})^{4}}\sqrt{\frac{24c^{4}c_{s}^{2}\kappa^{4}\mu^{2}r^{2}\log n}{m}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}.$
By triangle inequality and Lemmas 5.6 and 5.26 we have
$~{}\left\lvert\left\langle\eta\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2},\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{L}_{\natural}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\rangle\right\rvert$
$\leq$ $~{}\left\lvert\left\langle\eta\bm{L}_{\star}\bm{\Sigma}_{\star}^{-1/2}{\bm{R}}_{{\bm{A}}_{2}}^{*},\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\rangle\right\rvert$
$~{}+\left\lvert\left\langle\eta(\bm{L}_{\star}\bm{\Sigma}_{\star}^{-1/2}){\bm{R}}_{{\bm{A}}_{2}}^{*},\eta\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\right\rangle\right\rvert$
$\leq$ $~{}\varepsilon_{0}\eta^{2}\left\lVert{\bm{U}}_{\star}{\bm{R}}_{{\bm{A}}_{2}}^{*}\right\lVert_{\mathrm{F}}\left\lVert\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}$
$+\eta^{2}\sqrt{24p^{-1}n\log n}\left\lVert\bm{L}_{\star}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}_{2}}\right\lVert_{\mathrm{F}}$
$\leq$ $~{}\frac{\varepsilon_{0}\eta^{2}}{(1-\varepsilon_{0})^{2}}\left\lVert\bm{\Sigma}_{\star}^{1/2}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}^{2}+\frac{\eta^{2}(1+c\kappa)}{(1-\varepsilon_{0})^{3}}\sqrt{\frac{24c_{s}^{2}\mu^{2}r^{r}\log n}{m}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}},$
where in the third inequality follows from $\left\lVert{\bm{U}}_{\star}{\bm{R}}_{{\bm{A}}_{2}}^{*}\right\lVert_{\mathrm{F}}\leq\left\lVert{\bm{R}}_{{\bm{A}}_{2}}\right\lVert_{\mathrm{F}}$, $\left\lVert\bm{L}_{\star}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}=\left\lVert{\bm{U}}_{\star}\bm{\Sigma}_{\star}^{1/2}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}\leq\left\lVert\bm{\Sigma}_{\star}^{1/2}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}$ and (36), (30). Then, combing all the pieces, for $c\geq 1+\varepsilon_{0}$, $\varepsilon_{0},\eta\in(0,1)$, if provided $m\geq 24c^{2}(1+c\kappa)^{2}\varepsilon_{0}^{-2}c_{s}^{2}\mu^{2}r^{2}\log n$, we then have
$~{}\left\lvert\left\langle\bm{T}_{1},\bm{T}_{4}\right\rangle\right\rvert\leq\frac{(1-\eta)\eta\varepsilon_{0}}{1-\varepsilon_{0}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}+\frac{\varepsilon_{0}\eta^{2}}{(1-\varepsilon_{0})^{2}}\left\lVert\bm{\Sigma}_{\star}^{1/2}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}^{2}$
$\qquad\qquad\qquad+\frac{2\varepsilon_{0}\eta(1+\varepsilon_{0})}{(1-\varepsilon_{0})^{4}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}$
$\leq$ $\left(\frac{(1-\eta)\eta\varepsilon_{0}}{1-\varepsilon_{0}}+\frac{\varepsilon_{0}\eta(1+\varepsilon_{0})}{(1-\varepsilon_{0})^{4}}\right)\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}^{2}+\left(\frac{\varepsilon_{0}\eta^{2}}{(1-\varepsilon_{0})^{2}}+\frac{\varepsilon_{0}\eta(1+\varepsilon_{0})}{(1-\varepsilon_{0})^{4}}\right)\left\lVert\bm{\Sigma}_{\star}^{1/2}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\lVert_{\mathrm{F}}^{2}.$
For $\left\lVert\bm{T}_{4}\right\lVert_{\mathrm{F}}$: Using the variational representation of the Frobenius norm, for some ${\bm{L}}_{{\bm{A}}}\in\mathbb{C}^{n_{1}\times r}$, $\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}=1$ we have
$~{}\left\lVert\bm{T}_{4}\right\lVert_{\mathrm{F}}$
$=$ $~{}\eta\left\lvert\left\langle\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}+\bm{L}_{\natural}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right)\bm{R}_{\natural}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2},{\bm{L}}_{{\bm{A}}}\right\rangle\right\rvert$
$\leq$ $~{}\eta\left\lvert\left\langle\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}\right),{\bm{L}}_{{\bm{A}}}\bm{\Sigma}_{\star}^{1/2}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{R}_{\star}^{*}\right\rangle\right\rvert$
$+\eta\left\lvert\left\langle\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{\Delta}_{{\bm{L}}_{k}}\bm{R}_{\star}^{*}\right),{\bm{L}}_{{\bm{A}}}\bm{\Sigma}_{\star}^{1/2}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right\rangle\right\rvert$
$+\eta\left\lvert\left\langle\left(p^{-1}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}-\mathcal{G}\mathcal{G}^{*}\right)\left(\bm{L}_{\natural}\bm{\Delta}_{{\bm{R}}_{k}}^{*}\right),{\bm{L}}_{{\bm{A}}}\bm{\Sigma}_{\star}^{1/2}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{R}_{\natural}^{*}\right\rangle\right\rvert$
$\leq$ $~{}\varepsilon_{0}\eta\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}{\bm{V}}_{\star}\right\lVert_{\mathrm{F}}\left\lVert{\bm{L}}_{{\bm{A}}}\bm{\Sigma}_{\star}^{1/2}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}{\bm{V}}_{\star}\right\lVert_{\mathrm{F}}$
$+\eta\sqrt{24p^{-1}n\log n}\Big{(}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert\bm{R}_{\star}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}$
$+\left\lVert\bm{L}_{\natural}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\bm{\Sigma}_{\star}^{1/2}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}\left\lVert\bm{R}_{\star}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2,\infty}\Big{)}$
$\leq$ $~{}\frac{\varepsilon_{0}\eta}{(1-\varepsilon_{0})^{2}}\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}+\frac{2\eta(1+c\kappa)}{(1-\varepsilon_{0})^{3}}\sqrt{\frac{24c_{s}^{2}\mu^{2}r^{2}\log n}{m}}\left\lVert\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\right\lVert_{\mathrm{F}}$
where the second inequality follows from Lemmas 5.6 and 5.26, and the last inequality follows from (29c), (29d), (30), (36), as well as the following inequalities: $\|{\bm{L}}_{{\bm{A}}}\bm{\Sigma}_{\star}^{1/2}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}{\bm{V}}_{\star}^{*}\|_{\mathrm{F}}\leq\|{\bm{L}}_{{\bm{A}}}\bm{\Sigma}_{\star}^{1/2}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\|_{\mathrm{F}}\leq\frac{1}{(1-\varepsilon_{0})^{2}}$ and $\|\bm{\Delta}_{{\bm{R}}_{k}}\left(\bm{R}_{\natural}^{*}\bm{R}_{\natural}\right)^{-1}\bm{\Sigma}_{\star}^{1/2}\|_{\mathrm{F}}\leq\frac{1}{(1-\varepsilon_{0})^{2}}\|\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\|_{\mathrm{F}}\leq\frac{1}{\sigma^{\star}_{r}(1-\varepsilon_{0})^{2}}\|\bm{\Delta}_{{\bm{R}}_{k}}\bm{\Sigma}_{\star}^{1/2}\|_{\mathrm{F}}$.
6 Conclusion
In this paper, we introduce a novel non-convex algorithm, dubbed Hankel Structured Newton-Like Descent, for ill-conditioned robust Hankel recovery problems. The linear convergence guarantee has been established for the proposed algorithm; especially, its convergence rate is independent of the condition number of the underlying Hankel matrix. The empirical advantages of the proposed algorithm have been verified via extensive experiments on synthetic and real datasets.
Appendix A Proof of Lemma 5.26
Proof A.1.
First, we have
$~{}\left\lvert\left\langle\mathcal{G}\left(p^{-1}\Pi_{\Omega}-\mathcal{I}\right)\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right),{\bm{L}}_{{\bm{B}}}{\bm{R}}_{{\bm{B}}}^{*}\right\rangle\right\rvert$
$=$ $~{}p^{-1}\left\lvert\left\langle\left(\Pi_{\Omega}-p\mathcal{I}\right)\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right),\mathcal{G}^{*}\left({\bm{L}}_{{\bm{B}}}{\bm{R}}_{{\bm{B}}}^{*}\right)\right\rangle\right\rvert=p^{-1}\left\lvert\sum_{k=1}^{m}r_{k}\right\rvert,$
where
$r_{k}:=\left[\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right)\right]_{a_{k}}\left[\mathcal{G}^{*}\left({\bm{L}}_{{\bm{B}}}{\bm{R}}_{{\bm{B}}}^{*}\right)\right]_{a_{k}}-n^{-1}\left\langle\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right),\mathcal{G}^{*}\left({\bm{L}}_{{\bm{B}}}{\bm{R}}_{{\bm{B}}}^{*}\right)\right\rangle,~{}k\in[m],$
then we have $\mathbb{E}[r_{k}]=0$. Further, noticing that for any $a\in[n]$, we have
$~{}\left\lvert\left[\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right)\right]_{a}\left[\mathcal{G}^{*}\left({\bm{L}}_{{\bm{B}}}{\bm{R}}_{{\bm{B}}}^{*}\right)\right]_{a}\right\rvert$
$=$ $~{}\left\lvert\left(\sum_{(i,j)\in\Omega_{a}}\frac{1}{\varsigma_{a}}{\bm{L}}_{{\bm{A}}}(i,:){\bm{R}}_{{\bm{A}}}(j,:)^{*}\right)\left(\sum_{(i,j)\in\Omega_{a}}{\bm{L}}_{{\bm{B}}}(i,:){\bm{R}}_{{\bm{B}}}(j,:)^{*}\right)\right\rvert$
$\leq$ $~{}\left(\frac{1}{\varsigma_{a}}\sum_{(i,j)\in\Omega_{a}}\left\lVert{\bm{L}}_{{\bm{A}}}(i,:)\right\lVert_{2}\left\lVert{\bm{R}}_{{\bm{A}}}(j,:)\right\lVert_{2}\right)\left(\sum_{(i,j)\in\Omega_{a}}\left\lVert{\bm{L}}_{{\bm{B}}}(i,:)\right\lVert_{2}\left\lVert{\bm{R}}_{{\bm{B}}}(j,:)\right\lVert_{2}\right)$
$\leq$ $~{}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{B}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{B}}}\right\lVert_{\mathrm{F}},$
where the second and the inequalities follows from the Cauchy-Schwarz inequality and the fact that $\left\lvert\Omega_{a}\right\rvert=\varsigma_{a}$. Thus
$\left\lvert r_{k}\right\rvert\leq 2\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{B}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{B}}}\right\lVert_{\mathrm{F}}.$
Similarly, we have $\left\lvert r_{k}\right\rvert\leq 2\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{L}}_{{\bm{B}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{B}}}\right\lVert_{2,\infty}.$ Also, for any $a\in[n]$, by Cauchy-Schwarz inequality and $\left\lvert\Omega_{a}\right\rvert=\varsigma_{a}$ we have
$~{}\left\lvert\left[\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right)\right]_{a}\left[\mathcal{G}^{*}\left({\bm{L}}_{{\bm{B}}}{\bm{R}}_{{\bm{B}}}^{*}\right)\right]_{a}\right\rvert$
$\leq$ $~{}\left(\sum_{(i,j)\in\Omega_{a}}\frac{1}{\sqrt{\varsigma_{a}}}\left\lVert{\bm{L}}_{{\bm{A}}}(i,:)\right\lVert_{2}\left\lVert{\bm{R}}_{{\bm{A}}}(j,:)\right\lVert_{2}\right)\left(\frac{1}{\sqrt{\varsigma_{a}}}\sum_{(i,j)\in\Omega_{a}}\left\lVert{\bm{L}}_{{\bm{B}}}(i,:)\right\lVert_{2}\left\lVert{\bm{R}}_{{\bm{B}}}(j,:)\right\lVert_{2}\right)$
$\leq$ $\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{B}}}\right\lVert_{2,\infty}\left(\sum_{(i,j)\in\Omega_{a}}\frac{1}{\sqrt{\varsigma_{a}}}\left\lVert{\bm{R}}_{{\bm{A}}}(j,:)\right\lVert_{2}\right)\left(\frac{1}{\sqrt{\varsigma_{a}}}\sum_{(i,j)\in\Omega_{a}}\left\lVert{\bm{R}}_{{\bm{B}}}(j,:)\right\lVert_{2}\right)$
$\leq$ $\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{B}}}\right\lVert_{2,\infty}\sqrt{\sum_{i=1}^{a}\left\lVert{\bm{R}}_{{\bm{A}}}(i,:)\right\lVert_{2}^{2}}\sqrt{\sum_{j=1}^{a}\left\lVert{\bm{R}}_{{\bm{B}}}(j,:)\right\lVert_{2}^{2}}$
$\leq$ $\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{B}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{B}}}\right\lVert_{\mathrm{F}},$
Then we have $\left\lvert r_{k}\right\rvert\leq 2\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{B}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{B}}}\right\lVert_{\mathrm{F}}$. Similarly,
$\left\lvert r_{k}\right\rvert\leq 2\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{L}}_{{\bm{B}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{B}}}\right\lVert_{2,\infty}.$
Denote
$\begin{split}\zeta:=&~{}\min\{\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{B}}}\right\lVert_{\mathrm{F}},\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{L}}_{{\bm{B}}}\right\lVert_{2,\infty}\}\\ &~{}\times\min\{\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{B}}}\right\lVert_{\mathrm{F}},\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{B}}}\right\lVert_{2,\infty}\}\end{split}$
Then, combining all the pieces above we obtain
$\left\lvert r_{k}\right\rvert\leq 2\zeta.$
By using the above bounds for $\left\lvert\left[\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right)\right]_{a}\left[\mathcal{G}^{*}\left({\bm{L}}_{{\bm{B}}}{\bm{R}}_{{\bm{B}}}^{*}\right)\right]_{a}\right\rvert$ and the inequality $(x+y)^{2}\leq 2x^{2}+2y^{2}$, we further have
$\sum_{k=1}^{m}\mathbb{E}[r_{k}^{2}]\leq\sum_{k=1}^{m}4\max_{a\in[n]}\left\lvert\left[\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right)\right]_{a}\left[\mathcal{G}^{*}\left({\bm{L}}_{{\bm{B}}}{\bm{R}}_{{\bm{B}}}^{*}\right)\right]_{a}\right\rvert^{2}\leq 4m\zeta^{2}.$
By Bernstein’s inequality, it then implies
$\mathbb{P}\left(\left\lvert\sum_{k=1}^{m}r_{k}\right\rvert\geq t\right)\leq 2\exp\left(\frac{-t^{2}/2}{4m\zeta^{2}+\frac{4t}{3}\zeta}\right).$
Letting $t=\sqrt{24np\log n}\zeta$, by the definition of $r_{k}$ we then have
$\left\lvert\left\langle\mathcal{G}\left(p^{-1}\Pi_{\Omega}-\mathcal{I}\right)\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right),{\bm{L}}_{{\bm{B}}}{\bm{R}}_{{\bm{B}}}^{*}\right\rangle\right\rvert\leq p^{-1}t=\sqrt{24np^{-1}\log n}\zeta$
with probability at least $1-2n^{-2}$ provided $p\geq(24\log n)/n$.